### v2

# Order-Optimal Global Convergence for Average Reward Reinforcement Learning via Actor-Critic Approach

Abstract
This work analyzes average-reward reinforcement learning with general parametrization. Current state-of-the-art (SOTA) guarantees for this problem are either suboptimal or demand prior knowledge of the mixing time of the underlying Markov process, which is unavailable in most practical scenarios. We introduce a Multi-level Monte Carlo-based Natural Actor-Critic (MLMC-NAC) algorithm to address these issues. Our approach is the first to achieve a global convergence rate of $\tilde{\mathcal{O}}(1/\sqrt{T})$ without needing the knowledge of mixing time. It significantly surpasses the SOTA bound of $\tilde{\mathcal{O}}(T^{-1/4})$ where $T$ is the horizon length.
1 INTRODUCTION
Reinforcement Learning (RL) is a framework where an agent interacts with a Markovian environment and maximizes the reward it receives. The temporal dependence of the state transitions makes the problem of RL much more challenging than ordinary stochastic optimization, where data are selected in an independent and identically distributed (i.i.d.) manner. RL problems are typically analyzed via three setups: episodic, discounted reward with an infinite horizon, and average reward with an infinite horizon. The average reward framework is particularly significant for real-world applications, including robotics (Gonzalez et al., 2023), transportation (Al-Abbasi et al., 2019), communication networks (Agarwal et al., 2022), and healthcare (Tamboli et al., 2024).
One key approach to solving RL problems is model-based algorithms (Jaksch et al., 2010; Agrawal and Jia, 2017; Agarwal and Aggarwal, 2023) that directly learn the transition kernel from Markovian data. Unfortunately, such algorithms only work for small state spaces. Fueled by the practical need to learn policies for large state spaces, many model-free algorithms have emerged in the recent past where the policies are learned either from the $Q$-function or directly from the learned policy parameters. Some of these algorithms (Wei et al., 2020; Bai et al., 2024) use the observation that if the trajectory is divided into sub-trajectories that are mixing time apart, then these sub-trajectories become nearly independent. However, such approaches require the knowledge of mixing time which is not readily available in practice. Such an issue persists even if the MDP is assumed to be linear (Wei et al., 2020).
Recently, the authors of (Patel et al., 2024) showed that the algorithms that wait for orders of mixing time to have independent sub-trajectories require extremely large time horizons to even achieve sub-linear regret guarantees (even for mixing and hitting times of $10$, the time-horizon needed for sublinear guarantees is $6.6\times 10^{9}$). To alleviate these issues, (Patel et al., 2024) proposed an algorithm called Multi-level Actor Critic (MAC), where the estimates are obtained from Markovian trajectories utilizing the Multi-level Monte-Carlo (MLMC) gradient estimator (Giles, 2015) that does not require the knowledge of the mixing time. However, this approach is shown to achieve a global convergence rate of $\tilde{\operatorname*{\mathcal{O}}}(T^{-1/4})$ for general parameterized policies which is far away from the theoretical lower bound of $\Omega(T^{-1/2})$. This raises the following question.
Algorithm Mixing Time Global Model-Free? Policy
Unknown? Convergence Parametrization
UCRL2 (Jaksch et al., 2010) - ${\tilde{\operatorname*{\mathcal{O}}}(1/\sqrt{T})}$ No Tabular
PSRL (Agrawal and Jia, 2017) - ${\tilde{\operatorname*{\mathcal{O}}}(1/\sqrt{T})}$ No Tabular
FOPO (Wei et al., 2021) No ${\tilde{\operatorname*{\mathcal{O}}}(1/\sqrt{T})}$ No Linear
OLSVI.FH (Wei et al., 2021) No $\tilde{\operatorname*{\mathcal{O}}}(1/T^{1/4})$ No Linear
MDP-EXP2 (Wei et al., 2021) No ${\tilde{\operatorname*{\mathcal{O}}}(1/\sqrt{T})}$ No Linear
MDP-OOMD (Wei et al., 2020) No ${\tilde{\operatorname*{\mathcal{O}}}(1/\sqrt{T})}$ Yes Tabular
PPGAE (Bai et al., 2024) No $\tilde{\operatorname*{\mathcal{O}}}(1/T^{1/4})$ Yes General
NAC-CFA (Wang et al., 2024) No $\tilde{\operatorname*{\mathcal{O}}}(1/T^{1/4})$ Yes General
MAC (Patel et al., 2024) Yes $\tilde{\operatorname*{\mathcal{O}}}(1/T^{1/4})$ Yes General
MLMC-NAC (Algorithm 1) Yes ${\tilde{\operatorname*{\mathcal{O}}}(1/\sqrt{T})}$ Yes General
Table 1: Summary of the key results on global convergence guarantees for average reward reinforcement learning.
Is it possible to achieve a global convergence rate of $\tilde{\mathcal{O}}\left(T^{-1/2}\right)$ with general parameterized policies in the average-reward Markovian sampling setup without the algorithm requiring knowledge of the mixing time?
Our Contribution: This article answers this question affirmatively. In particular, we introduce a Multi-level Monte Carlo-based Natural Actor-Critic (MLMC-NAC) algorithm that comprises two major components. The first component, referred to as the Natural Policy Gradient (NPG) subroutine, obtains an approximate NPG direction which is used to update the policy parameter. One sub-task of obtaining the NPG is estimating the advantage function. This is done via the second component, known as the Critic subroutine that achieves its desired target via Temporal Difference (TD) learning. Both NPG and critic subroutines apply MLMC-based gradient estimators that eliminate the use of the mixing time in the algorithm. We establish that MLMC-NAC achieves a global convergence rate of $\tilde{\operatorname*{\mathcal{O}}}(T^{-1/2})$ which is optimal in the order of the horizon length $T$ and far surpasses the state-of-the-art rate of $\tilde{\operatorname*{\mathcal{O}}}\left(T^{-1/4}\right)$.
The primary difficulty in handling Markovian trajectories, as opposed to independent trajectories (Liu et al., 2020; Mondal and Aggarwal, 2024), is that an accurate estimation of the policy gradient often requires the knowledge of mixing and hitting times (Bai et al., 2024). In most practical cases, these quantities are unknown. Moreover, the hitting time scales with the size of the state space, limiting the application of the underlying algorithm. As stated earlier, similar to (Patel et al., 2024), we employ an MLMC-based actor-critic structure that allows us to establish guarantees for large or infinite state spaces while eliminating the use of mixing time. However, the improvement in our convergence result yields from the following novel introductions.
  • We use NPG instead of the PG method used in (Patel et al., 2024). We prove that the global convergence error can be bounded by the sum of terms proportional to the bias and second-order error in NPG estimation (Lemma 1).
  • Since the critic updates underlie the NPG subroutine, the NPG estimation errors are inherently connected to critic estimation errors. In particular, we exhibit that NPG bias and second-order error are determined by their critic counterparts (Theorem 3).
  • Finally, we bound the critic errors in Theorem 4 that, combined with Theorem 3 and Lemma 1, yields our desired convergence rate.
  • We observe that both NPG and critic routines can be viewed as stochastic linear recursions where the stochasticity arises due to MLMC-based estimation. Theorem 2 analyses the convergence rate of such a general recursion which is utilized to prove Theorem 3 and 4. We would like to emphasize that it is due to the careful analysis of convergence bias in Theorem 2 that we get the optimal rate of $\tilde{\operatorname*{\mathcal{O}}}(T^{-1/2})$. Had we performed a naive convergence bias analysis of the general stochastic linear recursion in Theorem 2, the convergence rate of our algorithm would have been far away from being optimal.
Related works: We will discuss the relevant works in three key areas as stated below. Our discussion primarily focuses on works that employ general parameterized policies. A summary of relevant works is available in Table 1.
  • Discounted Reward RL: We note that while optimal sample complexity has been studied for the discounted reward setups (Mondal and Aggarwal, 2024) for direct policy-gradient-based approaches, it assumes the availability of a simulator, enabling independent trajectory sampling. Markovian samplings are primarily studied via actor-critic approach (Xu et al., 2020; Gaur et al., 2024) which, unfortunately, have not achieved the optimal $\tilde{\operatorname*{\mathcal{O}}}(1/\epsilon^{2})$ sample complexity. Further, the analysis for discounted reward does not directly extend to the average reward setup since the sample complexity in the discounted case depends on $1/(1-\gamma)$, which blows up as $\gamma\to 1$.
  • Average Reward RL with direct Policy Gradient: The key work employing direct Policy Gradient for average reward setup with parametrized policy is (Bai et al., 2024). Unlike the discounted case, the policy gradient estimation is through the use of sub-trajectories that are large enough so that they are nearly independent. However, setting these trajectory lengths requires knowledge of the mixing time and the hitting time.
  • Average Reward RL with Actor-Critic approaches: To avoid creating independent subtrajectories, the policy gradient can be estimated by constructing value function and average reward estimates, updating based on observed samples. This is the basis of the actor-critic (AC) approach. Some studies on average reward reinforcement learning utilize an AC-based approach. For example, (Chen and Zhao, 2023) provides a local sample complexity of $\operatorname*{\mathcal{O}}(\epsilon^{-2})$ for vanilla AC, whereas (Panda and Bhatnagar, 2024) establishes a local sample complexity of $\mathcal{O}(\epsilon^{-2.08})$ for a Critic-Actor method. Another study, (Wang et al., 2024), introduces an AC method with a global sample complexity of $\mathcal{O}(\epsilon^{-4})$.
    Actor-critic methods employing Multi-Level Monte Carlo (MLMC)-based estimation were first explored in (Suttle et al., 2023) and subsequently in (Patel et al., 2024). The former study demonstrates a local convergence, while the latter establishes a global convergence rate. Both studies investigate an MLMC-based AC algorithm combined with AdaGrad (Duchi et al., 2011), inspired by stochastic gradient descent (SGD) related work in (Dorfman and Levy, 2022). Unfortunately, none of these studies lead to an optimal global convergence rate which is the goal of our work.
2 SETUP
In this paper, we explore an infinite horizon reinforcement learning problem with an average reward criterion, modeled by a Markov Decision Process (MDP) represented as a tuple $\mathcal{M}=(\mathcal{S},\mathcal{A},r,P,\rho)$. Here $\mathcal{S}$ indicates the state space, $\mathcal{A}$ defines the action space with a size of $A$, $r:\mathcal{S}\times\mathcal{A}\rightarrow[0,1]$ represents the reward function, $P:\mathcal{S}\times\mathcal{A}\rightarrow\Delta(\mathcal{S})$ defines the state transition function, where $\Delta(\mathcal{S})$ denotes the probability simplex over $\mathcal{S}$, and $\rho\in\Delta(\mathcal{S})$ signifies the initial distribution of states. A (stationary) policy $\pi:\mathcal{S}\rightarrow\Delta(\mathcal{A})$ determines the distribution of the action to be taken given the current state. It induces the following state transition $P^{\pi}:\mathcal{S}\rightarrow\Delta(\mathcal{S})$ given as $P^{\pi}(s,s^{\prime})=\sum_{a\in\mathcal{A}}P(s^{\prime}|s,a)\pi(a|s)$, $\forall s,s^{\prime}\in\mathcal{S}$. Observe that for any policy $\pi$, the sequence of states yielded by the MDP forms a Markov chain. We assume the following throughout the paper.
Assumption 1.
The Markov chain induced by every policy $\pi$, $\{s_{t}\}_{t\geq 0}$, is ergodic.
Before proceeding further, we point out that we consider a parameterized class of policies $\Pi$, which consists of all policies $\pi_{\theta}$ such that $\theta\in\Theta$, where $\Theta\subset\operatorname*{\mathbb{R}}^{\mathrm{d}}$. It is well-established that if $\mathcal{M}$ is ergodic, then $\forall\theta\in\Theta$, there exists a unique stationary $\rho$-independent distribution, denoted as $d^{\pi_{\theta}}\in\Delta(\mathcal{S})$, defined as follows.
$d^{\pi_{\theta}}(s)=\lim_{T\rightarrow\infty}\operatorname{\mathbb{E}}_{\pi_{\theta}}\left[\frac{1}{T}\sum_{t=0}^{T}\mathbf{1}(s_{t}=s)\bigg{|}s_{0}\sim\rho\right]$ (1)
where $\operatorname{\mathbb{E}}_{\pi_{\theta}}$ denotes the expectation over the distribution of all $\pi_{\theta}$-induced trajectories and $\mathbf{1}(\cdot)$ is an indicator function. The above distribution also obeys $P^{\pi_{\theta}}d^{\pi_{\theta}}=d^{\pi_{\theta}}$. With this notation in place, we define the mixing time of an MDP.
Definition 1.
The mixing time of an MDP $\mathcal{M}$ with respect to a policy parameter $\theta$ is defined as
$t_{\mathrm{mix}}^{\theta}\coloneqq\min\left\{t\geq 1\bigg{|}\|(P^{\pi_{\theta}})^{t}(s,\cdot)-d^{\pi_{\theta}}\|\leq\dfrac{1}{4},\forall s\in\mathcal{S}\right\}$
We define $t_{\mathrm{mix}}\coloneqq\sup_{\theta\in\Theta}t^{\theta}_{\mathrm{mix}}$ as the the overall mixing time. In this paper, $t_{\mathrm{mix}}$ is finite due to ergodicity.
The mixing time of an MDP measures how quickly the MDP approaches its stationary distribution when the same policy is executed repeatedly. In the average reward setting, we aim to find a policy $\pi_{\theta}$ that maximizes the long-term average reward defined below.
$J^{\pi_{\theta}}:=\lim_{T\rightarrow\infty}\mathbb{E}_{\pi_{\theta}}\left[\frac{1}{T}\sum_{t=0}^{T}r(s_{t},a_{t})\bigg{|}s_{0}\sim\rho\right]$ (2)
For simplicity, we denote $J(\theta)=J^{\pi_{\theta}}$. This paper uses an actor-critic approach to optimize $J$. Before proceeding further, we would like to introduce a few important terms. The action-value ($Q$) function corresponding to the policy $\pi_{\theta}$ is defined as
$Q^{\pi_{\theta}}(s,a)=$ $\mathbb{E}_{\pi_{\theta}}\Bigg{[}\sum_{t=0}^{\infty}\big{\{}r(s_{t},a_{t})-J(\theta)\big{\}}\bigg{|}s_{0}=s,a_{0}=a\Bigg{]}$
We can further define the state value function as
$V^{\pi_{\theta}}(s)=$ $\mathbb{E}_{a\sim\pi_{\theta}(\cdot|s)}[Q^{\pi_{\theta}}(s,a)]$ (3)
Bellman’s equation, thus, takes the following form (Puterman, 2014)
$Q^{\pi_{\theta}}(s,a)$ $=r(s,a)-J(\theta)+\mathbb{E}[V^{\pi_{\theta}}(s^{\prime})],$ (4)
where the expectation is over $s^{\prime}\sim P(\cdot|s,a)$. We define the advantage as $A^{\pi_{\theta}}(s,a)\triangleq Q^{\pi_{\theta}}(s,a)-V^{\pi_{\theta}}(s).$ With the notations in place, we express below the well-known policy gradient theorem established by (Sutton et al., 1999).
$\nabla_{\theta}J(\theta)\!=\!\mathbb{E}_{(s,a)\sim\nu^{\pi_{\theta}}}\!\bigg{[}\!A^{\pi_{\theta}}(s,a)\nabla_{\theta}\log\pi_{\theta}(a|s)\!\bigg{]}$ (5)
where $\nu^{\pi_{\theta}}(s,a)=d^{\pi_{\theta}}(s)\pi(a|s)$. Policy Gradient (PG) algorithms maximize the average reward by updating $\theta$ along the policy gradient $\nabla_{\theta}J(\theta)$. In contrast, Natural Policy Gradient (NPG) methods update $\theta$ along the NPG direction $\omega^{*}_{\theta}$ where
$\omega^{*}_{\theta}=F(\theta)^{\dagger}\nabla_{\theta}J(\theta),$ (6)
The symbol $\dagger$ denotes the Moore-Penrose pseudoinverse and $F(\theta)$ is the Fisher matrix as defined as
$F(\theta)=\operatorname{\mathbb{E}}_{(s,a)\sim\nu^{\pi_{\theta}}}\left[\nabla_{\theta}\log\pi_{\theta}(a|s)\otimes\nabla_{\theta}\log\pi_{\theta}(a|s)\right]$ (7)
where $\otimes$ symbolizes the outer product. The precoder $F(\theta)$ takes the change of the parameterized policy with respect to $\theta$ into account, thereby preventing overshooting or slow updates of $\theta$. Note that $\omega_{\theta}^{*}$ can be written as the minimizer of the function $L_{\nu^{\pi_{\theta}}}(\cdot,\theta)$ where
$L_{\nu^{\pi_{\theta}}}(\omega,\theta)\textstyle{=\dfrac{1}{2}\operatorname{\mathbb{E}}_{(s,a)\sim\nu^{\pi_{\theta}}}\left[\big{(}A^{\pi_{\theta}}(s,a)-\omega^{\top}\nabla_{\theta}\log\pi_{\theta}(a|s)\big{)}^{2}\right]}$ (8)
for all $\omega\in\mathbb{R}^{\mathrm{d}}$. This is essentially a convex optimization that can be iteratively solved utilizing a gradient-based method. Invoking (5), one can show that
$\nabla_{\omega}L_{\nu^{\pi_{\theta}}}(\omega,\theta)=F(\theta)\omega-\nabla_{\theta}J(\theta)$ (9)
Note that $\nabla_{\omega}L_{\nu^{\pi_{\theta}}}(\omega,\theta)$ is not exactly computable since the transition function $P$ and hence the stationary distribution, $d^{\pi_{\theta}}$, and the advantage function, $A^{\pi_{\theta}}(\cdot,\cdot)$ are typically unknown in most practical cases. Recall that $A^{\pi_{\theta}}(s,a)=Q^{\pi_{\theta}}(s,a)-V^{\pi_{\theta}}(s)$. Moreover, Bellman’s equation (4) states that $Q^{\pi_{\theta}}(s,a)$ is determined by $J(\theta)$ and $V^{\pi_{\theta}}$. Notice that $J(\theta)$ can be written as a solution to the following optimization problem.
$\min_{\eta\in\mathbb{R}}R(\theta,\eta)\coloneqq\frac{1}{2}\sum_{s\in\mathcal{S}}\sum_{a\in\mathcal{A}}\nu^{\pi_{\theta}}(s,a)\left\{\eta-r(s,a)\right\}^{2}$ (10)
Algorithm 1 Multi-level Monte Carlo-based Natural Actor-Critic (MLMC-NAC)
1:Input: Initial parameters $\theta_{0}$, $\{\omega_{H}^{k}\}$, and $\{\xi_{0}^{k}\}$, policy update stepsize $\alpha$, parameters for NPG update, $\gamma$, parameters for critic update, $\beta$, $c_{\beta}$, initial state $s_{0}\sim\rho$, outer loop size $K$, inner loop size $H$, $T_{\max}$
2:Initialization: $T\leftarrow 0$
3:for $k=0,1,\cdots,K-1$ do
4:     $\triangleright$ Average reward and critic estimation
5:     for $h=0,1,\cdots,H-1$ do
6:         $s_{0}^{kh}\leftarrow s_{0}$, $Q_{kh}\sim\text{Geom}(1/2)$
7:         for $t=0,\dots,\max\{2^{Q_{kh}},T_{\max}\}-1$ do
8:              Take action $a^{kh}_{t}\sim\pi_{\theta_{k}}(\cdot|s^{kh}_{t})$
9:              Collect next state $s^{kh}_{t+1}\sim P(\cdot|s_{t}^{kh},a^{kh}_{t})$
10:              Receive reward $r(s_{t}^{kh},a_{t}^{kh})$
11:         end for
12:         $T_{kh}\leftarrow\max\{2^{Q_{kh}},T_{\max}\}$, $s_{0}\leftarrow s^{kh}_{T_{kh}}$
13:         Update $\xi_{h}^{k}$ using (15) and (3)
14:         $T\leftarrow T+T_{kh}$
15:     end for
16:     $\xi_{k}\leftarrow\xi_{H}^{k}$
17:     $\triangleright$ Natural Policy Gradient (NPG) estimation
18:     for $h=H,H+1,\cdots,2H-1$ do
19:         $s_{0}^{kh}\leftarrow s_{0}$, $Q_{kh}\sim\text{Geom}(1/2)$
20:         for $t=0,\dots,\max\{2^{Q_{kh}},T_{\max}\}-1$ do
21:              Take action $a_{t}^{kh}\sim\pi_{\theta_{k}}(\cdot|s_{t}^{kh})$
22:              Collect next state $s^{kh}_{t+1}\sim P(\cdot|s_{t}^{kh},a_{t}^{kh})$
23:              Receive reward $r(s_{t}^{kh},a_{t}^{kh})$
24:         end for
25:         $T_{kh}\leftarrow\max\{2^{Q_{kh}},T_{\max}\}$, $s_{0}\leftarrow s^{kh}_{T_{kh}}$
26:         Update $\omega_{h}^{k}$ using (16) and (20)
27:         $T\leftarrow T+T_{kh}$
28:     end for
29:     $\triangleright$ Policy update
30:     $\omega_{k}\leftarrow\omega^{k}_{2H}$, $\theta_{k+1}\leftarrow\theta_{k}+\alpha\omega_{k}$
31:end for
The above formulation allows us to compute $J(\theta)$ in a gradient-based iterative manner. In particular,
$\nabla_{\eta}R(\theta,\eta)=\sum_{s\in\mathcal{S}}\sum_{a\in\mathcal{A}}\nu^{\pi_{\theta}}(s,a)\left\{\eta-r(s,a)\right\}$ (11)
To facilitate the estimation of the advantage function, we assume that $V^{\pi_{\theta}}(\cdot)$ can be approximated by a critic function $\hat{V}(\zeta_{\theta},\cdot)=(\zeta_{\theta})^{\top}\phi(\cdot)$ where $\zeta_{\theta}\in\mathbb{R}^{m}$ denotes a solution to the following optimization problem, and $\phi(s)\in\mathbb{R}^{m}$, $\left\lVert\phi(s)\right\rVert\leq 1$ is a feature vector, $\forall s\in\mathcal{S}$.
$\min_{\zeta\in\mathbb{R}^{m}}E(\theta,\zeta)\coloneqq\frac{1}{2}\sum_{s\in\mathcal{S}}d^{\pi_{\theta}}(s)(V^{\pi_{\theta}}(s)-\hat{V}(\zeta,s))^{2}$ (12)
The above formulation paves a way to compute $\zeta_{\theta}$ via a gradient-based iterative procedure. Note the following.
$\begin{split}&\nabla_{\zeta}E(\theta,\zeta)=\sum_{s\in\mathcal{S}}\sum_{a\in\mathcal{A}}\nu^{\pi_{\theta}}(s,a)\left\{\zeta^{\top}\phi(s)-Q^{\pi_{\theta}}(s,a)\right\}\phi(s)\end{split}$ (13)
The iterative updates of $\theta$, $\eta$, and $\zeta$, along their associated gradients provided in (9), (11), and (LABEL:eq_nu_grad) form the basis of our algorithm stated next.
3 PROPOSED ALGORITHM
We propose a Multi-level Monte Carlo-based Natural Actor-Critic (MLMC-NAC) algorithm (Algorithm 1) that runs $K$ number of epochs (also called outer loops). The $k$th loop obtains $\xi_{k}=[\eta_{k},\zeta_{k}]^{\top}$ where $\eta_{k}$ denotes an estimate of the average reward $J(\theta_{k})$, and $\zeta_{k}$ defines an estimate of the critic parameter, $\zeta_{\theta_{k}}$. These estimates are then used to compute the approximate NPG, $\omega_{k}$, which is applied to update the policy parameter $\theta_{k}$.
$\theta_{k+1}=\theta_{k}+\alpha\omega_{k}$ (14)
where $\alpha$ is the learning parameter. The estimate $\xi_{k}$ is obtained in $H$ number of inner loop steps. In particular, $\forall h\in\{0,\cdots,H-1\}$, we apply the following updates starting from arbitrary $\xi_{0}^{k}$, and finally assign $\xi_{H}^{k}=\xi_{k}$.
$\begin{split}&\xi^{k}_{h+1}=\xi^{k}_{h}-\beta\mathbf{v}_{h}(\theta_{k},\xi^{k}_{h}),~{}\text{where}~{}\mathbf{v}_{h}(\theta_{k},\xi^{k}_{h})=\left[c_{\beta}\hat{\nabla}_{\eta}R(\theta_{k},\eta^{k}_{h}),\hat{\nabla}_{\zeta}E(\theta_{k},\xi^{k}_{h})\right]^{\top}\end{split}$ (15)
where $c_{\beta}$ is a constant, and $\beta$ is a learning rate. Observe that ^$\hat{\nabla}_{\zeta}E(\theta_{k},\xi_{h}^{k})$ has dependence on $\xi_{h}^{k}=[\eta_{h}^{k},\zeta_{h}^{k}]^{\top}$ due to the presence of $Q$-function in (LABEL:eq_nu_grad) while ^$\hat{\nabla}_{\eta}R(\theta_{k},\eta_{h}^{k})$ is dependent only on $\eta_{h}^{k}$ (details given later).
The approximate NPG $\omega_{k}$ is also obtained in $H$ inner loop steps, starting from arbitrary $\omega^{k}_{H}$, then recursively applying the stochastic gradient descent (SGD) method as stated below $\forall h\in\{H,\cdots,2H-1\}$, and assigning $\omega_{k}=\omega^{k}_{2H}$ [^1] We initiate from $h=H$, instead of $h=0$ to emphasize that the iteration (16) starts after $H$ iterations of (15)..
$\begin{split}\omega^{k}_{h+1}&=\omega^{k}_{h}-\gamma\hat{\nabla}_{\omega}L_{\nu^{\pi_{\theta_{k}}}}(\omega^{k}_{h},\theta_{k},\xi_{k})\end{split}$ (16)
where $\gamma$ is a learning parameter, and ^$\hat{\nabla}_{\omega}L_{\nu^{\pi_{\theta}}}(\omega^{k}_{h},\theta_{k},\xi_{k})$ symbolizes an estimate of $\nabla_{\omega}L_{\nu^{\pi_{\theta_{k}}}}(\omega^{k}_{h},\theta_{k})$. We would like to clarify that the above gradient estimate depends on $\xi_{k}$ because of the presence of the advantage function in the expression of the policy gradient whose estimation needs both $\eta_{k}$, $\zeta_{k}$ (details given later). A process is stated below to calculate the gradient estimates used in (16) and (15).
Gradient Estimation via MLMC: Consider a $\pi_{\theta_{k}}$-induced trajectory $\mathcal{T}_{kh}=\{(s_{t}^{kh},a_{t}^{kh},s_{t+1}^{kh})\}_{t=0}^{l_{kh}-1}$ whose length is given as $l_{kh}=2^{Q_{kh}}$ where $Q_{kh}\sim\mathrm{Geom}(\frac{1}{2})$. We can write the $Q$ estimate as below $\forall t\in\{0,\cdots,l_{kh}-1\}$ using Bellman’s equation (4).
$\hat{Q}^{\pi_{\theta_{k}}}(\xi_{k},z_{t}^{kh})=r(s_{t}^{kh},a_{t}^{kh})-\eta_{k}+\zeta_{k}^{\top}\phi(s_{t+1}^{kh})$ (17)
where $z_{t}^{kh}\coloneqq(s_{t}^{kh},a^{kh}_{t},s^{kh}_{t+1})$. This leads to the advantage estimate (also called the temporal difference error) as follows $\forall t\in\{0,\cdots,l_{kh}-1\}$.
$\begin{split}\hat{A}^{\pi_{\theta_{k}}}(\xi_{k},z_{t}^{kh})&=r(s_{t}^{kh},a_{t}^{kh})-\eta_{k}+\zeta_{k}^{\top}\left[\phi(s_{t+1}^{kh})-\phi(s_{t}^{kh})\right]\end{split}$ (18)
Define the following quantity $\forall t\in\{0,\cdots,l_{kh}-1\}$.
$\mathbf{u}_{t}^{kh}=\hat{A}_{\mathbf{u}}(\theta_{k},z_{t}^{kh})\omega_{h}^{k}-\hat{b}_{\mathbf{u}}(\theta_{k},\xi_{k},z_{t}^{kh})~{}\text{where}$ (19)
$\hat{A}_{\mathbf{u}}(\theta_{k},z_{t}^{kh})=\nabla_{\theta}\log\pi_{\theta_{k}}(a_{t}^{kh}|s_{t}^{kh})\otimes\nabla_{\theta}\log\pi_{\theta_{k}}(a_{t}^{kh}|s_{t}^{kh})$
$\hat{b}_{\mathbf{u}}(\theta_{k},\xi_{k},z_{t}^{kh})=\hat{A}^{\pi_{\theta_{k}}}(\xi_{k},z_{t}^{kh})\nabla_{\theta}\log\pi_{\theta_{k}}(a_{t}^{kh}|s_{t}^{kh})$
The term $\mathbf{u}^{kh}_{t}$ can be thought of as a crude estimate of $\nabla_{\omega}L_{\nu^{\pi_{\theta_{k}}}}(\omega^{k}_{h},\theta_{k})$, obtained using a single transition $z_{t}^{kh}$. One naive way to refine this estimate is to calculate an empirical average of $\{\mathbf{u}^{kh}_{t}\}_{t=0}^{l_{kh}-1}$. In this work, however, we resort to the MLMC method where the final estimate is given as follows.
^$\begin{split}\hat{\nabla}_{\omega}L_{\nu^{\pi_{\theta_{k}}}}(\omega^{k}_{h},\theta_{k},\xi_{k})&=U_{0}+\begin{cases}2^{Q}\left(U^{Q}-U^{Q-1}\right),&\text{if}~{}2^{Q}\leq T_{\max}\\ 0&\text{otherwise}\end{cases}\\ &=\hat{A}^{\mathrm{MLMC}}_{\mathbf{u},k,h}(\theta_{k})\omega_{h}^{k}-\hat{b}^{\mathrm{MLMC}}_{\mathbf{u},k,h}(\theta_{k},\xi_{k})\end{split}$ (20)
where $U^{j}=\frac{1}{2^{j}}\sum_{t=1}^{2^{j}}\mathbf{u}_{t}^{kh}$, $j\in\{Q-1,Q\}$, $T_{\max}\geq 2$ is a parameter, and $\hat{A}^{\mathrm{MLMC}}_{\mathbf{u},k,h}(\theta_{k})$, and $\hat{b}^{\mathrm{MLMC}}_{\mathbf{u},k,h}(\theta_{k},\xi_{k})$ denote MLMC-based estimates of samples $\{\hat{A}_{\mathbf{u}}(\theta_{k},z_{t}^{kh})\}_{t=0}^{l_{kh}-1}$ and $\{\hat{b}_{\mathbf{u}}(\theta_{k},\xi_{k},z_{t}^{kh})\}_{t=0}^{l_{kh}-1}$ respectively.
The advantage of MLMC is that it generates the same order of bias as the empirical average of $T_{\max}$ samples but requires only $\mathcal{O}(\log T_{\max})$ samples on an average. It also eliminates the requirement for the knowledge of the mixing time which was crucial in some earlier works (Bai et al., 2024) that employed empirical average-based estimates.
Using a similar method, we will now obtain an estimate of $\mathbf{v}_{h}(\theta_{k},\xi^{k}_{h})$. Following our earlier notations, we denote $\mathcal{T}_{kh}=\{(s^{kh}_{t},a^{kh}_{t})\}_{t=0}^{l_{kh}-1}$ as a $\pi_{\theta_{k}}$-induced trajectory of length $l_{kh}=2^{Q_{kh}}$, where $Q_{kh}\sim\mathrm{Geom}(\frac{1}{2})$. Notice the terms stated below $\forall t\in\{0,\cdots,l_{kh}-1\}$.
$\begin{split}\mathbf{v}^{kh}_{t}&=\begin{bmatrix}c_{\beta}\left\{\eta_{h}^{k}-r(s_{t}^{kh},a_{t}^{kh})\right\}\\ \bigg{\{}(\zeta_{h}^{k})^{\top}\phi(s_{t}^{kh})-\hat{Q}^{\pi_{\theta_{k}}}(\xi_{h}^{k},z_{t}^{kh})\bigg{\}}\phi(s_{t}^{kh})\end{bmatrix}=\hat{A}_{\mathbf{v}}(z_{t}^{kh})\xi^{k}_{h}-\hat{b}_{\mathbf{v}}(z_{t}^{kh})\end{split}$ (21)
where $z_{t}^{kh}\coloneqq(s_{t}^{kh},a_{t}^{kh},s^{kh}_{t+1})$, $\hat{Q}^{\pi_{\theta_{k}}}(\xi_{h}^{k},z_{t}^{kh})$ is given by (17) and $\hat{A}_{\mathbf{v}}(z_{t}^{kh})$, $\hat{b}_{\mathbf{v}}(z_{t}^{kh})$ are defined as
$\hat{A}_{\mathbf{v}}(z_{t}^{kh})$ $=\begin{bmatrix}c_{\beta}&0\\ \phi(s_{t}^{kh})&\phi(s_{t}^{kh})\left[\phi(s_{t}^{kh})-\phi(s_{t+1}^{kh})\right]^{\top}\end{bmatrix},$ (22)
$\hat{b}_{\mathbf{v}}(z_{t}^{kh})$ $=\begin{bmatrix}c_{\beta}r(s_{t}^{kh},a_{t}^{kh})\\ r(s_{t}^{kh},a_{t}^{kh})\phi(s_{t}^{kh})\end{bmatrix}$ (23)
Based on (11) and (LABEL:eq_nu_grad), the term $\mathbf{v}^{kh}_{t}$ can be thought of as a crude approximation of $\mathbf{v}_{h}(\theta_{k},\xi_{h}^{k})$ obtained using a single transition, $z_{t}^{kh}$. The final estimate is
$\mathbf{v}_{h}(\theta_{k},\xi^{k}_{h})$ $=V_{0}+\begin{cases}2^{Q}\left(V^{Q}-V^{Q-1}\right),&\text{if}~{}2^{Q}\leq T_{\max}\\ 0&\text{otherwise}\end{cases}$
$=\hat{A}_{\mathbf{v},k,h}^{\mathrm{MLMC}}\xi_{h}^{k}-\hat{b}_{\mathbf{v},k,h}^{\mathrm{MLMC}}$ (24)
where $V^{j}\coloneqq 2^{-j}\sum_{t=1}^{2^{j}}\mathbf{v}_{t}^{kh}$, $j\in\{Q-1,Q\}$. Moreover, $\hat{A}_{\mathbf{v},k,h}^{\mathrm{MLMC}}$ and $\hat{b}_{\mathbf{v},k,h}^{\mathrm{MLMC}}$ symbolize MLMC-based estimates of $\{\hat{A}_{\mathbf{v}}(z_{t}^{kh})\}_{t=0}^{l_{kh}-1}$ and $\{\hat{b}_{\mathbf{v}}(z_{t}^{kh})\}_{t=0}^{l_{kh}-1}$ respectively.
A few remarks are in order. Although the MLMC-based estimates achieve the same order of bias as the empirical average with a lower average sample requirement, its variance is larger. Many existing literature reduce the impact of the increased variance via AdaGrad-based parameter updates. Though such methods typically work well for general non-convex optimization problems, it does not exploit any inherent structure of strongly convex optimization problems, thereby being sub-optimal for both the NPG-finding subroutine and the average reward and critic updates. In this paper, we resort to a version of the standard SGD to cater to our following needs. First, by judiciously choosing the learning parameters, we prove that it is possible to achieve the optimal rate without invoking AdaGrad-type updates. Moreover, the choice of our parameters does not require the knowledge of the mixing time. Finally, although our gradient estimates suffer from bias due to the inherent error present in the critic approximation, our novel analysis suitably handles these issues.
4 MAIN RESULTS
We first state some relevant assumptions. Let $A_{\mathbf{v}}(\theta)\coloneqq\operatorname{\mathbb{E}}_{\theta}\left[\hat{A}_{\mathbf{v}}(z)\right]$, and $b_{\mathbf{v}}(\theta)\coloneqq\operatorname{\mathbb{E}}_{\theta}\left[\hat{b}_{\mathbf{v}}(z)\right]$ where matrices $\hat{A}_{\mathbf{v}}(z)$, $\hat{b}_{\mathbf{v}}(z)$ are described in (22), (23) and the expectation $\operatorname{\mathbb{E}}_{\theta}$ is computed over the distribution of $z=(s,a,s^{\prime})$ where $(s,a)\sim\nu^{\pi_{\theta}}$, $s^{\prime}\sim P(\cdot|s,a)$. For arbitrary policy parameter $\theta$, we denote $\xi_{\theta}^{*}=[A_{\mathbf{v}}(\theta)]^{\dagger}b_{\mathbf{v}}(\theta)=[\eta_{\theta}^{*},\zeta_{\theta}^{*}]^{\top}$. Using these notations, below we state some assumptions related to the critic analysis.
Assumption 2.
We assume the critic approximation error, $\epsilon_{\mathrm{app}}$ (defined below) is finite.
$\epsilon_{\mathrm{app}}=\sup_{\theta}~{}E(\theta,\zeta_{\theta}^{*})$ (25)
Assumption 3.
There exist $\lambda>0$ such that $\forall\theta$
$\operatorname{\mathbb{E}}_{\theta}[\phi(s)(\phi(s)-\phi(s^{\prime}))^{\top}]\succcurlyeq\lambda I$ (26)
[…]
Both Assumptions 2 and 3 are frequently employed in the analysis of actor-critic methods (Suttle et al., 2023; Patel et al., 2024; Wu et al., 2020; Panda and Bhatnagar, 2024). Assumption 2 intuitively relates to the quality of the feature mapping where $\epsilon_{\mathrm{app}}$ measures the quality. Well-designed feature maps may lead to small or even zero $\epsilon_{\mathrm{app}}$, whereas poorly designed features result in higher errors. Assumption 3 is essential for guaranteeing the uniqueness of the solution to the minimization problem (12). Assumption 3 also follows when the set of policy parameters, $\Theta$ is compact and $e\notin W_{\phi}$, where $e$ is the vector of all ones and $W_{\phi}$ is the space spanned by the feature vectors. To see this, note that if $e\notin W_{\phi}$, there exists $\lambda_{\theta}$ for every policy $\pi_{\theta}$ such that $\operatorname{\mathbb{E}}_{\theta}[\phi(s)(\phi(s)-\phi(s^{\prime}))^{\top}]\succcurlyeq\lambda_{\theta}I$ (Zhang et al., 2021b). Since $\Theta$ is compact, setting $\lambda=\inf_{\theta\in\Theta}\lambda_{\theta}>0$ satisfies Assumption 3.
For large enough $c_{\beta}$, Assumption 3 implies that $A_{\mathbf{v}}(\theta)-(\lambda/2)I$ is positive definite (refer to Lemma 8). It also implies that $A_{\mathbf{v}}(\theta)$ is invertible. We will now state some assumptions related to the policy parameterization.
Assumption 4.
For any $\theta$, the transferred compatible function approximation error, $L_{\nu^{\pi^{*}}}(\omega_{\theta}^{*};\theta)$, satisfies the following inequality.
$L_{d^{\pi^{*}}}(\omega_{\theta}^{*};\theta)\leq\epsilon_{\mathrm{bias}}$
where $\omega_{\theta}^{*}$ denotes the exact NPG direction at $\theta$ defined by (6), $\pi^{*}$ indicates the optimal policy, and the function $L_{\nu^{\pi^{*}}}(\cdot,\cdot)$ is given by (8).
Assumption 5.
For all $\theta,\theta_{1},\theta_{2}$ and $(s,a)\in\mathcal{S}\times\mathcal{A}$, the following statements hold.
$(a)\text{ }\|\nabla_{\theta}\log\pi_{\theta}(a|s)\|\leq G_{1}$
$(b)\text{ }\|\nabla_{\theta}\log\pi_{\theta_{1}}(a|s)-\nabla_{\theta}\log\pi_{\theta_{2}}(a|s)\|\leq G_{2}\|\theta_{1}-\theta_{2}\|$
Assumption 6 (Fisher non-degenerate policy).
There exists a constant $\mu>0$ such that $F(\theta)-\mu I_{d}$ is positive semidefinite where $I_{d}$ denotes an identity matrix.
Comments on Assumptions 4-6: We would like to highlight that all these assumptions are commonly found in PG literature (Liu et al., 2020; Agarwal et al., 2021; Papini et al., 2018; Xu et al., 2019; Fatkhullin et al., 2023). We elaborate more on these assumptions below.
The term $\epsilon_{\mathrm{bias}}$ captures the expressivity of the parameterized policy class. If, for example, the policy class is complete such as in the case of softmax parametrization, $\epsilon_{\mathrm{bias}}=0$ (Agarwal et al., 2021). However, for restricted parametrization which may not contain all stochastic policies, we have $\epsilon_{\mathrm{bias}}>0$. It is known that $\epsilon_{\mathrm{bias}}$ is insignificant for rich neural parametrization (Wang et al., 2019). Note that Assumption 5 requires that the score function is bounded and Lipschitz continuous. This assumption is widely used in the analysis of PG-based methods (Liu et al., 2020; Agarwal et al., 2021; Papini et al., 2018; Xu et al., 2019; Fatkhullin et al., 2023). Assumption 6 requires that the eigenvalues of the Fisher information matrix can be bounded from below and is commonly used in obtaining global complexity bounds for PG-based procedures (Liu et al., 2020; Zhang et al., 2021a; Bai et al., 2022; Fatkhullin et al., 2023). Assumptions 5-6 were shown to hold for various examples recently including Gaussian policies with linearly parameterized means and certain neural parametrizations (Liu et al., 2020; Fatkhullin et al., 2023).
With the relevant assumptions in place, we are now ready to state our main result.
Theorem 1.
Consider Algorithm 1 with $K=\Theta(\sqrt{T})$, $H=\Theta(\sqrt{T}/\log(T))$. Let Assumptions 1-6 hold and $J$ be $L$-smooth. There exists a choice of parameters such that the following holds for a sufficiently large $T$.
$J^{*}-\frac{1}{K}\sum_{k=0}^{K-1}\operatorname{\mathbb{E}}[J(\theta_{k})]$ $\leq\operatorname*{\mathcal{O}}\left(\sqrt{\epsilon_{\mathrm{app}}}+\sqrt{\epsilon_{\mathrm{bias}}}\right)+\tilde{\operatorname*{\mathcal{O}}}\left(t_{\mathrm{mix}}^{3}T^{-1/2}\right)$ (27)
where $J^{*}$ is the optimal value of $J(\cdot)$.
The values of the learning parameters used in the above theorem can be found in Appendix H. It is to be mentioned that the above bound of $\tilde{\operatorname*{\mathcal{O}}}(1/\sqrt{T})$ is a significant improvement in comparison to the state-of-the-art bounds of $\tilde{\operatorname*{\mathcal{O}}}(1/T^{1/4})$ in the average reward general parameterization setting (Bai et al., 2024; Patel et al., 2024). Also, our bounds do not depend on the size of the action space and hitting time unlike that in (Bai et al., 2024). Although (Patel et al., 2024) provides bounds with $\operatorname*{\mathcal{O}}(\sqrt{t_{\mathrm{mix}}})$ dependence, these bounds, unfortunately, depend on the projection radius of the critic updates, $R_{\omega}$, which can be large and scale with $t_{\mathrm{mix}}$ (Wei et al., 2020). In contrast, our algorithm does not use such projection operators and therefore, does not scale with $R_{\omega}$.
5 PROOF OUTLINE
We segregate our analysis into four parts as described below.
5.1 Policy update analysis
Lemma 1.
Consider any policy update rule of form
$\theta_{k+1}=\theta_{k}+\alpha\omega_{k}.$ (28)
If Assumptions 4 and 5 hold, then the following inequality is satisfied for any $K$.
$J^{*}-\frac{1}{K}\sum_{k=0}^{K-1}\operatorname{\mathbb{E}}[J(\theta_{k})]\leq\sqrt{\epsilon_{\mathrm{bias}}}$ $+\frac{G_{1}}{K}\sum_{k=0}^{K-1}\operatorname{\mathbb{E}}\|(\operatorname{\mathbb{E}}_{k}\left[\omega_{k}\right]-\omega^{*}_{k})\|+\frac{\alpha G_{2}}{2K}\sum_{k=0}^{K-1}\operatorname{\mathbb{E}}\|\omega_{k}\|^{2}$
$+\frac{1}{\alpha K}\operatorname{\mathbb{E}}_{s\sim d^{\pi^{*}}}[KL(\pi^{*}(\cdot|s)\|\pi_{\theta_{0}}(\cdot|s))]$ (29)
where $KL(\cdot\|\cdot)$ is the Kullback-Leibler divergence, $\omega^{*}_{k}$ is the NPG direction $F(\theta_{k})^{-1}\nabla J(\theta_{k})$, $\pi^{*}$ is the optimal policy, $J^{*}$ is the optimal value of the function $J(\cdot)$, and $\operatorname{\mathbb{E}}_{k}[\cdot]$ denotes conditional expectation given the history up to epoch $k$.
Note that the last term in (1) is $\operatorname*{\mathcal{O}}(1/K)$. The term $\operatorname{\mathbb{E}}\|\omega_{k}\|^{2}$ can be further decomposed as
$\begin{split}\operatorname{\mathbb{E}}\|\omega_{k}\|^{2}&\leq 2\operatorname{\mathbb{E}}\|\omega_{k}-\omega_{k}^{*}\|^{2}+2\operatorname{\mathbb{E}}\|\omega_{k}^{*}\|^{2}\overset{(a)}{\leq}2\operatorname{\mathbb{E}}\|\omega_{k}-\omega_{k}^{*}\|^{2}+\dfrac{2}{\mu^{2}}\operatorname{\mathbb{E}}\|\nabla_{\theta}J(\theta_{k})\|^{2}\end{split}$ (30)
where $(a)$ follows from Assumption 6 and the definition that $\omega_{k}^{*}=F(\theta_{k})^{-1}\nabla_{\theta}J(\theta_{k})$. Further, it can be proven that for the choice of $\alpha$ used in Theorem 1, we have
$\begin{split}\dfrac{1}{\mu^{2}K}&\left(\sum_{k=0}^{K-1}\|\nabla_{\theta}J(\theta_{k})\|^{2}\right)\leq\dfrac{32LG_{1}^{4}}{\mu^{4}K}+\left(\dfrac{2G_{1}^{4}}{\mu^{2}}+1\right)\left(\dfrac{1}{K}\sum_{k=0}^{K-1}\|\omega_{k}-\omega_{k}^{*}\|^{2}\right)\end{split}$ (31)
Evidently, one can obtain a global convergence bound by bounding the terms $\operatorname{\mathbb{E}}\|\omega_{k}-\omega^{*}_{k}\|^{2}$, and $\operatorname{\mathbb{E}}\|(\operatorname{\mathbb{E}}_{k}\left[\omega_{k}\right]-\omega^{*}_{k})\|$. These terms define the second-order error and bias of the NPG estimator, $\omega_{k}$. In the next subsections, we briefly describe how to obtain these bounds.
5.2 Analysis of a General Linear Recursion
Observe that the NPG finding subroutine (16) and the update of the critic parameter and the average reward (15) can be written in the following form for a given $k$.
$x_{h+1}=x_{h}-\bar{\beta}(\hat{P}_{h}x_{h}-\hat{q}_{h})$ (32)
where $\hat{P}_{h}$, $\hat{q}_{h}$ are MLMC based estimates of the matrices $P\in\mathbb{R}^{n\times n}$, $q\in\mathbb{R}^{n}$ respectively, and $h\in\{0,\cdots,H-1\}$. Assume that the following bounds hold $\forall h$.
$\begin{split}&\operatorname{\mathbb{E}}_{h}\left[\left\lVert\hat{P}_{h}-P\right\rVert^{2}\right]\leq\sigma_{P}^{2},~{}\left\lVert\operatorname{\mathbb{E}}_{h}\left[\hat{P}_{h}\right]-P\right\rVert^{2}\leq\delta^{2}_{P},\\ &\operatorname{\mathbb{E}}_{h}\left[\left\lVert\hat{q}_{h}-q\right\rVert^{2}\right]\leq\sigma_{q}^{2},~{}\left\lVert\operatorname{\mathbb{E}}_{h}\left[\hat{q}_{h}\right]-q\right\rVert^{2}\leq\delta_{q}^{2},~{}\text{and~{}}\left\lVert\operatorname{\mathbb{E}}\left[\hat{q}_{h}\right]-q\right\rVert^{2}\leq\bar{\delta}_{q}^{2}\end{split}$ (33)
where $\operatorname{\mathbb{E}}_{h}$ denotes conditional expectation given history up to step $h$. Since $\operatorname{\mathbb{E}}[\hat{q}_{h}]=\operatorname{\mathbb{E}}[\operatorname{\mathbb{E}}_{h}[\hat{q}_{h}]]$, we have $\bar{\delta}^{2}_{q}\leq\delta_{q}^{2}$. Additionally, assume that
$\begin{split}0<\lambda_{P}\leq\left\lVert P\right\rVert\leq\Lambda_{P}~{}~{}\text{and}~{}\left\lVert q\right\rVert\leq\Lambda_{q}\end{split}$ (34)
The condition that $\lambda_{P}>0$ implies that $P$ is invertible. The goal of recursion (32) is to approximate the term $x^{*}=P^{-1}q$. We have the following result.
Theorem 2.
Consider the recursion (32). Assume that the conditions (33), and (34) hold. Also, let $\delta_{P}\leq\lambda_{P}/8$, and $\bar{\beta}=\frac{2\log H}{\lambda_{P}H}$. The following relation holds whenever $H$ is sufficiently large.
$\operatorname{\mathbb{E}}\left[\|x_{H}-x^{*}\|^{2}\right]\leq\dfrac{\operatorname{\mathbb{E}}\left[\|x_{0}-x^{*}\|^{2}\right]}{H^{2}}+\tilde{\operatorname*{\mathcal{O}}}\bigg{(}\dfrac{R_{0}}{H}+R_{1}\bigg{)}$
where $R_{0}=\lambda_{P}^{-4}\Lambda_{q}^{2}\sigma_{P}^{2}+\lambda_{P}^{-2}\sigma_{q}^{2}$, $R_{1}=\lambda_{P}^{-2}\big{[}\delta_{P}^{2}\lambda_{P}^{-2}\Lambda_{q}^{2}+\delta_{q}^{2}\big{]}$, and $\tilde{\operatorname*{\mathcal{O}}}(\cdot)$ hides logarithmic factors of $H$. Moreover,
$\|\operatorname{\mathbb{E}}[x_{H}]-x^{*}\|^{2}\leq\frac{\|\operatorname{\mathbb{E}}[x_{0}]-x^{*}\|^{2}}{H^{2}}+\operatorname*{\mathcal{O}}(\bar{R}_{1})+\operatorname*{\mathcal{O}}\left(\lambda_{P}^{-2}\delta_{P}^{2}\left\{\operatorname{\mathbb{E}}\left[\|x_{0}-x^{*}\|^{2}\right]+\mathcal{O}\left(R_{0}+R_{1}\right)\right\}\right)$
where $\bar{R}_{1}=\lambda_{P}^{-2}\big{[}\delta_{P}^{2}\lambda_{P}^{-2}\Lambda_{q}^{2}+\bar{\delta}_{q}^{2}\big{]}$.
We shall now use Theorem 2 to characterize the estimation errors in the NPG-finding subroutine and average reward and critic updates.
5.3 Analysis of NPG-Finding Subroutine
In the NPG finding subroutine, the goal is to compute $\omega_{k}^{*}=[F(\theta_{k})]^{-1}\nabla_{\theta}J(\theta_{k})$. An estimate of $F(\theta_{k})$ is given by $\hat{A}^{\mathrm{MLMC}}_{\mathbf{u},k,h}(\theta_{k})$, and that of the policy gradient $\nabla_{\theta}J(\theta_{k})$ is given by $\hat{b}^{\mathrm{MLMC}}_{\mathbf{u},k,h}(\theta_{k},\xi_{k})$ (see (20)). One can establish the following inequalities invoking the properties of the MLMC estimates.
Lemma 2.
Fix an instant $k$ of the outer loop in Algorithm 1. Given $(\theta_{k},\xi_{k})$, the MLMC estimates defined in (20) satisfy the following bounds $\forall h\in\{H,\cdots,2H-1\}$ provided the assumptions in Theorem 1 hold.
$(a)~{}\left\lVert\operatorname{\mathbb{E}}_{k,h}\left[\hat{A}^{\mathrm{MLMC}}_{\mathbf{u},k,h}(\theta_{k})\right]-F(\theta_{k})\right\rVert^{2}\leq\operatorname*{\mathcal{O}}\left(G_{1}^{4}t_{\mathrm{mix}}T_{\max}^{-1}\right)$
$(b)~{}\operatorname{\mathbb{E}}_{k,h}\left[\left\lVert\hat{A}^{\mathrm{MLMC}}_{\mathbf{u},k,h}(\theta_{k})-F(\theta_{k})\right\rVert^{2}\right]\leq\operatorname*{\mathcal{O}}\left(G_{1}^{4}t_{\mathrm{mix}}\log T_{\max}\right)$
$(c)~{}\left\lVert\operatorname{\mathbb{E}}_{k,h}\left[\hat{b}^{\mathrm{MLMC}}_{\mathbf{u},k,h}(\theta_{k},\xi_{k})\right]-\nabla_{\theta}J(\theta_{k})\right\rVert^{2}\leq\operatorname*{\mathcal{O}}\left(\sigma_{\mathbf{u},k}^{2}t_{\mathrm{mix}}T_{\max}^{-1}+\delta_{\mathbf{u},k}^{2}\right)$
$(d)~{}\operatorname{\mathbb{E}}_{k,h}\left[\left\lVert\hat{b}^{\mathrm{MLMC}}_{\mathbf{u},k,h}(\theta_{k},\xi_{k})-\nabla_{\theta}J(\theta_{k})\right\rVert^{2}\right]\leq\operatorname*{\mathcal{O}}\left(\sigma_{\mathbf{u},k}^{2}t_{\mathrm{mix}}\log T_{\max}+\delta_{\mathbf{u},k}^{2}\right)$
where $\operatorname{\mathbb{E}}_{k,h}$ defines the conditional expectation given the history up to the inner loop step $h$ (within the $k$th outer loop instant), $\sigma_{\mathbf{u},k}^{2}=\operatorname*{\mathcal{O}}\left(G_{1}^{2}\left\lVert\xi_{k}\right\rVert^{2}\right)$ and
$\delta_{\mathbf{u},k}^{2}=\mathcal{O}\left(G_{1}^{2}\left\lVert\xi_{k}-\xi_{k}^{*}\right\rVert^{2}+G_{1}^{2}\epsilon_{\mathrm{app}}\right)$
where $\xi_{k}^{*}\coloneqq\xi_{\theta_{k}}^{*}=[A_{\mathbf{v}}(\theta_{k})]^{-1}b_{\mathbf{v}}(\theta_{k})$ and $\operatorname{\mathbb{E}}_{k}$ defines the conditional expectation given the history up to the outer loop instant $k$. Moreover, given $\theta_{k}$, we also have
$\begin{split}&(e)~{}\left\lVert\operatorname{\mathbb{E}}_{k}\left[\hat{b}^{\mathrm{MLMC}}_{\mathbf{u},k,h}(\theta_{k},\xi_{k})\right]-\nabla_{\theta}J(\theta_{k})\right\rVert^{2}\leq\operatorname*{\mathcal{O}}\left(\bar{\sigma}_{\mathbf{u},k}^{2}t_{\mathrm{mix}}T_{\max}^{-1}+\bar{\delta}_{\mathbf{u},k}^{2}\right)\end{split}$
where $\bar{\sigma}^{2}_{\mathbf{u},k}=\operatorname*{\mathcal{O}}(G_{1}^{2}\operatorname{\mathbb{E}}_{k}\left\lVert\xi_{k}\right\rVert^{2})$, and $\bar{\delta}^{2}_{\mathbf{u},k}$ is given as
$\bar{\delta}_{\mathbf{u},k}^{2}=\mathcal{O}\left(G_{1}^{2}\left\lVert\operatorname{\mathbb{E}}_{k}[\xi_{k}]-\xi_{k}^{*}\right\rVert^{2}+G_{1}^{2}\epsilon_{\mathrm{app}}\right)$
Combining Lemma 2 and Theorem 2, we arrive at the following results.
Theorem 3.
Consider the NPG-finding recursion (16) with $\gamma=\frac{2\log H}{\mu H}$ and $T_{\max}=H^{2}$. If all assumptions in Theorem 1 hold, then for sufficiently large $c_{\beta},H$
$\mathbb{E}_{k}\left[\|\omega_{k}-\omega_{k}^{*}\|^{2}\right]\leq\frac{1}{H^{2}}\|\omega^{k}_{H}-\omega_{k}^{*}\|^{2}+\tilde{\operatorname*{\mathcal{O}}}\left(\frac{G_{1}^{6}t_{\mathrm{mix}}^{3}}{\mu^{4}H}+\dfrac{G_{1}^{2}c_{\beta}^{2}t_{\mathrm{mix}}}{\mu^{2}\lambda^{2}H}\right)+\mu^{-2}G_{1}^{2}\operatorname*{\mathcal{O}}\big{(}\operatorname{\mathbb{E}}_{k}\|\xi_{k}-\xi_{k}^{*}\|^{2}+\epsilon_{\mathrm{app}}\big{)}$
Additionally, we also have
$\|\mathbb{E}_{k}[\omega_{k}]-\omega_{k}^{*}\|^{2}$ $\leq\tilde{\operatorname*{\mathcal{O}}}\left(\dfrac{G_{1}^{4}t_{\mathrm{mix}}}{\mu^{2}H^{2}}\left\lVert\omega_{H}^{k}-\omega_{k}^{*}\right\rVert^{2}\right)+\mu^{-2}G_{1}^{2}\operatorname*{\mathcal{O}}\bigg{(}\|\operatorname{\mathbb{E}}_{k}[\xi_{k}]-\xi_{k}^{*}\|^{2}+\epsilon_{\mathrm{app}}\bigg{)}$
$+\tilde{\operatorname*{\mathcal{O}}}\left(\dfrac{G_{1}^{6}t^{2}_{\mathrm{mix}}}{\mu^{4}H^{2}}\operatorname{\mathbb{E}}_{k}\left[\left\lVert\xi_{k}-\xi_{k}^{*}\right\rVert^{2}\right]\right)+\tilde{\operatorname*{\mathcal{O}}}\left(\dfrac{G_{1}^{4}t_{\mathrm{mix}}}{\mu^{2}H^{2}}\left\{\mu^{-4}G_{1}^{6}t_{\mathrm{mix}}^{3}+\mu^{-2}\lambda^{-2}G_{1}^{2}c_{\beta}^{2}t_{\mathrm{mix}}\right\}\right)$
5.4 Critic and Average Reward Analysis
The goal of the recursion (15) is to compute the term $\xi_{k}^{*}=[A_{\mathbf{v}}(\theta_{k})]^{-1}b_{\mathbf{v}}(\theta_{k})$. An estimate of $A_{\mathbf{v}}(\theta_{k})$ is given by $\hat{A}_{\mathbf{v},k,h}^{\mathrm{MLMC}}$ while that of $b_{\mathbf{v}}(\theta_{k})$ is given by $\hat{b}_{\mathbf{v},k,h}^{\mathrm{MLMC}}$ (see (3)). Similar to Lemma 2, we have the following result.
Lemma 3.
Given the parameter $\theta_{k}$, the MLMC estimates defined in (3) obey the following bounds provided the assumptions in Theorem 1 hold.
$(a)~{}\left\lVert\operatorname{\mathbb{E}}_{k,h}\left[\hat{A}^{\mathrm{MLMC}}_{\mathbf{v},k,h}\right]-A_{\mathbf{v}}(\theta_{k})\right\rVert^{2}\leq\operatorname*{\mathcal{O}}\left(c_{\beta}^{2}t_{\mathrm{mix}}T_{\max}^{-1}\right)$
$(b)~{}\operatorname{\mathbb{E}}_{k,h}\left[\left\lVert\hat{A}^{\mathrm{MLMC}}_{\mathbf{v},k,h}-A_{\mathbf{v}}(\theta_{k})\right\rVert^{2}\right]\leq\operatorname*{\mathcal{O}}\left(c_{\beta}^{2}t_{\mathrm{mix}}\log T_{\max}\right)$
$(c)~{}\left\lVert\operatorname{\mathbb{E}}_{k,h}\left[\hat{b}^{\mathrm{MLMC}}_{\mathbf{v},k,h}\right]-b_{\mathbf{v}}(\theta_{k})\right\rVert^{2}\leq\operatorname*{\mathcal{O}}\left(c_{\beta}^{2}t_{\mathrm{mix}}T_{\max}^{-1}\right)$
$(d)~{}\operatorname{\mathbb{E}}_{k,h}\left[\left\lVert\hat{b}^{\mathrm{MLMC}}_{\mathbf{v},k,h}-b_{\mathbf{v}}(\theta_{k})\right\rVert^{2}\right]\leq\operatorname*{\mathcal{O}}\left(c_{\beta}^{2}t_{\mathrm{mix}}\log T_{\max}\right)$
where $h\in\{0,1,\cdots,H-1\}$ and $\operatorname{\mathbb{E}}_{k,h}$ is interpreted in the same way as in Lemma 2.
Lemma 3 and Theorem 2 lead to the following.
Theorem 4.
Consider the recursion (3). Let $T_{\max}=H^{2}$, $\beta=\frac{4\log H}{\lambda H}$. If all assumptions of Theorem 1 hold, then the following is true for sufficiently large $c_{\beta},H$.
$\mathbb{E}_{k}\left[\|\xi_{k}-\xi_{k}^{*}\|^{2}\right]\leq\frac{1}{H^{2}}\|\xi^{k}_{0}-\xi_{k}^{*}\|^{2}+\tilde{\operatorname*{\mathcal{O}}}\left(\frac{c_{\beta}^{4}t_{\mathrm{mix}}}{\lambda^{4}H}\right),$
$\|\operatorname{\mathbb{E}}_{k}[\xi_{k}]-\xi_{k}^{*}\|^{2}\leq\operatorname*{\mathcal{O}}\left(\dfrac{c_{\beta}^{2}t_{\mathrm{mix}}}{\lambda^{2}H^{2}}\left\lVert\xi_{0}^{k}-\xi_{k}^{*}\right\rVert^{2}+\dfrac{c_{\beta}^{6}t^{2}_{\mathrm{mix}}}{\lambda^{6}H^{2}}\right)$
Combining Lemma 1, Theorem 3 and 4, we establish Theorem 1. We would like to emphasize the importance of the term $\bar{\delta}_{q}^{2}$ in Theorem 2. A naive analysis would have resulted in a worse upper bound in Theorem 2 that replaces $\bar{\delta}_{q}^{2}$ with $\delta_{q}^{2}$. Such degradation in Theorem 2 would have resulted in a convergence rate of $\tilde{\operatorname*{\mathcal{O}}}(T^{-1/3})$ as opposed to our current bound of $\tilde{\operatorname*{\mathcal{O}}}(T^{-1/2})$. Finally, it is to be mentioned that our convergence bound does not depend on $|\mathcal{S}|$, thereby enabling its application to large state space MDPs as long as $t_{\mathrm{mix}}$ is finite.
6 CONCLUSIONS
This work presents the Multi-Level Monte Carlo-based Natural Actor-Critic (MLMC-NAC) algorithm for addressing average-reward reinforcement learning challenges. The proposed method achieves an order-optimal global convergence rate of $\tilde{\mathcal{O}}(1/\sqrt{T})$, significantly surpassing the state-of-the-art results in this domain, particularly for model-free approaches with general policy parametrization without the knowledge of the mixing time. Future work includes extensions of our ideas to constrained setups, general utility RL, etc.
Appendix A Proof of Lemma 1
The proof of the lemma is straightforward and follows a similar structure to the one in [Mondal and Aggarwal, 2024]. The major distinction is that the bound derived in [Mondal and Aggarwal, 2024] applies to the discounted reward setting, whereas our derivation pertains to the average-reward case. We begin by stating a useful lemma.
Lemma 4 (Lemma 4, [Bai et al., 2024]).
The difference in the performance for any policies $\pi_{\theta}$ and $\pi_{\theta^{\prime}}$ is bounded as follows
$J(\theta)-J(\theta^{\prime})=\operatorname{\mathbb{E}}_{s\sim d^{\pi_{\theta}}}\operatorname{\mathbb{E}}_{a\sim\pi_{\theta}(\cdot|s)}\big{[}A^{\pi_{\theta^{\prime}}}(s,a)\big{]}.$ (35)
Continuing with the proof, we have:
$\operatorname{\mathbb{E}}_{s\sim d^{\pi^{*}}}[KL(\pi^{*}(\cdot|s)\|\pi_{\theta_{k}}(\cdot|s))-KL(\pi^{*}(\cdot|s)\|\pi_{\theta_{k+1}}(\cdot|s))]$
$=\operatorname{\mathbb{E}}_{s\sim d^{\pi^{*}}}\operatorname{\mathbb{E}}_{a\sim\pi^{*}(\cdot|s)}\bigg{[}\log\frac{\pi_{\theta_{k+1}(a|s)}}{\pi_{\theta_{k}}(a|s)}\bigg{]}$
$\overset{(a)}{\geq}\operatorname{\mathbb{E}}_{s\sim d^{\pi^{*}}}\operatorname{\mathbb{E}}_{a\sim\pi^{*}(\cdot|s)}[\nabla_{\theta}\log\pi_{\theta_{k}}(a|s)\cdot(\theta_{k+1}-\theta_{k})]-\frac{G_{2}}{2}\|\theta_{k+1}-\theta_{k}\|^{2}$
$=\alpha\operatorname{\mathbb{E}}_{s\sim d^{\pi^{*}}}\operatorname{\mathbb{E}}_{a\sim\pi^{*}(\cdot|s)}[\nabla_{\theta}\log\pi_{\theta_{k}}(a|s)\cdot\omega_{k}]-\frac{G_{2}\alpha^{2}}{2}\|\omega_{k}\|^{2}$
$=\alpha\operatorname{\mathbb{E}}_{s\sim d^{\pi^{*}}}\operatorname{\mathbb{E}}_{a\sim\pi^{*}(\cdot|s)}[\nabla_{\theta}\log\pi_{\theta_{k}}(a|s)\cdot\omega^{*}_{k}]+\alpha\operatorname{\mathbb{E}}_{s\sim d^{\pi^{*}}}\operatorname{\mathbb{E}}_{a\sim\pi^{*}(\cdot|s)}[\nabla_{\theta}\log\pi_{\theta_{k}}(a|s)\cdot(\omega_{k}-\omega^{*}_{k})]-\frac{G_{2}\alpha^{2}}{2}\|\omega_{k}\|^{2}$
$=\alpha[J^{*}-J(\theta_{k})]+\alpha\operatorname{\mathbb{E}}_{s\sim d^{\pi^{*}}}\operatorname{\mathbb{E}}_{a\sim\pi^{*}(\cdot|s)}[\nabla_{\theta}\log\pi_{\theta_{k}}(a|s)\cdot\omega^{*}_{k}]-\alpha[J^{*}-J(\theta_{k})]$
$\hskip 28.45274pt+\alpha\operatorname{\mathbb{E}}_{s\sim d^{\pi^{*}}}\operatorname{\mathbb{E}}_{a\sim\pi^{*}(\cdot|s)}[\nabla_{\theta}\log\pi_{\theta_{k}}(a|s)\cdot(\omega_{k}-\omega^{*}_{k})]-\frac{G_{2}\alpha^{2}}{2}\|\omega_{k}\|^{2}$
=$\overset{(b)}{=}\alpha[J^{*}-J(\theta_{k})]+\alpha\operatorname{\mathbb{E}}_{s\sim d^{\pi^{*}}}\operatorname{\mathbb{E}}_{a\sim\pi^{*}(\cdot|s)}\bigg{[}\nabla_{\theta}\log\pi_{\theta_{k}}(a|s)\cdot\omega^{*}_{k}-A^{\pi_{\theta_{k}}}(s,a)\bigg{]}$
$\hskip 28.45274pt+\alpha\operatorname{\mathbb{E}}_{s\sim d^{\pi^{*}}}\operatorname{\mathbb{E}}_{a\sim\pi^{*}(\cdot|s)}[\nabla_{\theta}\log\pi_{\theta_{k}}(a|s)\cdot(\omega_{k}-\omega^{*}_{k})]-\frac{G_{2}\alpha^{2}}{2}\|\omega_{k}\|^{2}$
$\overset{(c)}{\geq}\alpha[J^{*}-J(\theta_{k})]-\alpha\sqrt{\operatorname{\mathbb{E}}_{s\sim d^{\pi^{*}}}\operatorname{\mathbb{E}}_{a\sim\pi^{*}(\cdot|s)}\bigg{[}\bigg{(}\nabla_{\theta}\log\pi_{\theta_{k}}(a|s)\cdot\omega^{*}_{k}-A^{\pi_{\theta_{k}}}(s,a)\bigg{)}^{2}\bigg{]}}$
$\hskip 28.45274pt+\alpha\operatorname{\mathbb{E}}_{s\sim d^{\pi^{*}}}\operatorname{\mathbb{E}}_{a\sim\pi^{*}(\cdot|s)}[\nabla_{\theta}\log\pi_{\theta_{k}}(a|s)\cdot(\omega_{k}-\omega^{*}_{k})]-\frac{G_{2}\alpha^{2}}{2}\|\omega_{k}\|^{2}$
$\overset{(d)}{\geq}\alpha[J^{*}-J(\theta_{k})]-\alpha\sqrt{\epsilon_{\mathrm{bias}}}+\alpha\operatorname{\mathbb{E}}_{s\sim d^{\pi^{*}}}\operatorname{\mathbb{E}}_{a\sim\pi^{*}(\cdot|s)}[\nabla_{\theta}\log\pi_{\theta_{k}}(a|s)\cdot(\omega_{k}-\omega^{*}_{k})]-\frac{G_{2}\alpha^{2}}{2}\|\omega_{k}\|^{2}.$
Here $(a)$ and $(b)$ follow from Assumption 5$(b)$ and Lemma 4, respectively. Inequality $(c)$ results from the convexity of the function $f(x)=x^{2}$. Lastly, $(d)$ is a consequence of Assumption 4. By taking expectations on both sides, we derive:
$\begin{split}&\operatorname{\mathbb{E}}\left[\operatorname{\mathbb{E}}_{s\sim d^{\pi^{*}}}\left[KL(\pi^{*}(\cdot|s)\|\pi_{\theta_{k}}(\cdot|s))-KL(\pi^{*}(\cdot|s)\|\pi_{\theta_{k+1}}(\cdot|s))\right]\right]\\ &\geq\alpha[J^{*}-\operatorname{\mathbb{E}}\left[J(\theta_{k})]\right]-\alpha\sqrt{\epsilon_{\mathrm{bias}}}\\ &\hskip 56.9055pt+\alpha\operatorname{\mathbb{E}}\left[\operatorname{\mathbb{E}}_{s\sim d^{\pi^{*}}}\operatorname{\mathbb{E}}_{a\sim\pi^{*}(\cdot|s)}[\nabla_{\theta}\log\pi_{\theta_{k}}(a|s)\cdot(\operatorname{\mathbb{E}}_{k}[\omega_{k}]-\omega^{*}_{k})]\right]-\frac{G_{2}\alpha^{2}}{2}\operatorname{\mathbb{E}}\left[\|\omega_{k}\|^{2}\right]\\ &\overset{}{\geq}\alpha[J^{*}-\operatorname{\mathbb{E}}\left[J(\theta_{k})]\right]-\alpha\sqrt{\epsilon_{\mathrm{bias}}}\\ &\hskip 56.9055pt-\alpha\operatorname{\mathbb{E}}\left[\operatorname{\mathbb{E}}_{s\sim d^{\pi^{*}}}\operatorname{\mathbb{E}}_{a\sim\pi^{*}(\cdot|s)}[\|\nabla_{\theta}\log\pi_{\theta_{k}}(a|s)\|\|\operatorname{\mathbb{E}}_{k}[\omega_{k}]-\omega^{*}_{k}\|]\right]-\frac{G_{2}\alpha^{2}}{2}\operatorname{\mathbb{E}}\left[\|\omega_{k}\|^{2}\right]\\ &\overset{(a)}{\geq}\alpha[J^{*}-\operatorname{\mathbb{E}}\left[J(\theta_{k})]\right]-\alpha\sqrt{\epsilon_{\mathrm{bias}}}-\alpha G_{1}\operatorname{\mathbb{E}}\|(\operatorname{\mathbb{E}}_{k}[\omega_{k}]-\omega^{*}_{k})\|-\frac{G_{2}\alpha^{2}}{2}\operatorname{\mathbb{E}}\left[\|\omega_{k}\|^{2}\right]\end{split}$ (36)
where $(a)$ follows from Assumption 5$(a)$. Rearranging the terms, we get,
$\begin{split}J^{*}-\operatorname{\mathbb{E}}[J(\theta_{k})]&\leq\sqrt{\epsilon_{\mathrm{bias}}}+G_{1}\operatorname{\mathbb{E}}\|(\operatorname{\mathbb{E}}_{k}[\omega_{k}]-\omega^{*}_{k})\|+\frac{G_{2}\alpha}{2}\operatorname{\mathbb{E}}\|\omega_{k}\|^{2}\\ &+\frac{1}{\alpha}\operatorname{\mathbb{E}}\left[\operatorname{\mathbb{E}}_{s\sim d^{\pi^{*}}}[KL(\pi^{*}(\cdot|s)\|\pi_{\theta_{k}}(\cdot|s))-KL(\pi^{*}(\cdot|s)\|\pi_{\theta_{k+1}}(\cdot|s))]\right]\end{split}$ (37)
Adding the above inequality from $k=0$ to $K-1$, using the non-negativity of KL divergence, and dividing the resulting expression by $K$, we obtain the final result.
Appendix B Proof of Theorem 2
Let $g_{h}=\hat{P}_{h}x_{h}-\hat{q}_{h}$. To prove the first statement, observe the following relations.
$\|x_{h+1}-x^{*}\|^{2}$ $=\|x_{h}-\bar{\beta}{g}_{h}-x^{*}\|^{2}$
$=\|x_{h}-x^{*}\|^{2}-2\bar{\beta}\langle x_{h}-x^{*},{g}_{h}\rangle+\bar{\beta}^{2}\|g_{h}\|^{2}$
$\overset{}{=}\|x_{h}-x^{*}\|^{2}-2\bar{\beta}\langle x_{h}-x^{*},P(x_{h}-x^{*})\rangle-2\bar{\beta}\langle x_{h}-x^{*},{g}_{h}-P(x_{h}-x^{*})\rangle+\bar{\beta}^{2}\|g_{h}\|^{2}$
$\overset{(a)}{\leq}\|x_{h}-x^{*}\|^{2}-2\bar{\beta}\lambda_{P}\|x_{h}-x^{*}\|^{2}-2\bar{\beta}\langle x_{h}-x^{*},{g}_{h}-P(x_{h}-x^{*})\rangle$
$\hskip 28.45274pt+2\bar{\beta}^{2}\|g_{h}-P(x_{h}-x^{*})\|^{2}+2\bar{\beta}^{2}\|P(x_{h}-x^{*})\|^{2}$
$\overset{(b)}{\leq}\|x_{h}-x^{*}\|^{2}-2\bar{\beta}\lambda_{P}\|x_{h}-x^{*}\|^{2}-2\bar{\beta}\langle x_{h}-x^{*},g_{h}-P(x_{h}-x^{*})\rangle$
$\hskip 28.45274pt+2\bar{\beta}^{2}\|g_{h}-P(x_{h}-x^{*})\|^{2}+2\Lambda_{P}^{2}\bar{\beta}^{2}\|x_{h}-x^{*}\|^{2}$
where $(a)$ and $(b)$ follow from $\lambda_{P}\leq\left\lVert P\right\rVert\leq\Lambda_{P}$. Taking conditional expectation $\operatorname{\mathbb{E}}_{h}$ on both sides, we obtain
$\operatorname{\mathbb{E}}_{h}\left[\left\|x_{h+1}-x^{*}\right\|^{2}\right]$ $\leq(1-2\bar{\beta}\lambda_{P}+2\Lambda_{P}^{2}\bar{\beta}^{2})\|x_{h}-x^{*}\|^{2}-2\bar{\beta}\langle x_{h}-x^{*},\operatorname{\mathbb{E}}_{h}\left[g_{h}-P(x_{h}-x^{*})\right]\rangle$
$\hskip 28.45274pt+2\bar{\beta}^{2}\operatorname{\mathbb{E}}_{h}\left\|g_{h}-P(x_{h}-x^{*})\right\|^{2}$ (38)
Observe that the third term in (B) can be bounded as follows.
$\|g_{h}-P(x_{h}-x^{*})\|^{2}$ $=\|(\hat{P}_{h}-P)(x_{h}-x^{*})+(\hat{P}_{h}-P)x^{*}+(q-\hat{q}_{h})\|^{2}$
$\leq 3\|\hat{P}_{h}-P\|^{2}\|x_{h}-x^{*}\|^{2}+3\|\hat{P}_{h}-P\|^{2}\|x^{*}\|^{2}+3\|q-\hat{q}_{h}\|^{2}$
$\leq 3\|\hat{P}_{h}-P\|^{2}\|x_{h}-x^{*}\|^{2}+3\lambda_{P}^{-2}\Lambda_{q}^{2}\|\hat{P}_{h}-P\|^{2}+3\|q-\hat{q}_{h}\|^{2}$
where the last inequality follows from $\left\lVert x^{*}\right\rVert^{2}=\left\lVert P^{-1}q\right\rVert^{2}\leq\lambda_{P}^{-2}\Lambda_{q}^{2}$. Taking expectation yields
$\operatorname{\mathbb{E}}_{h}\|g_{h}-P(x_{h}-x^{*})\|^{2}$ $\leq 3\operatorname{\mathbb{E}}_{h}\|\hat{P}_{h}-P\|^{2}\|x_{h}-x^{*}\|^{2}+3\lambda_{P}^{-2}\Lambda_{q}^{2}\operatorname{\mathbb{E}}_{h}\|\hat{P}_{h}-P\|^{2}+3\operatorname{\mathbb{E}}_{h}\|\hat{q}_{h}-q\|^{2}$
$\leq 3\sigma_{P}^{2}\left\lVert x_{h}-x^{*}\right\rVert^{2}+3\lambda_{P}^{-2}\Lambda_{q}^{2}\sigma_{P}^{2}+3\sigma_{q}^{2}$ (39)
The second term in (B) can be bounded as
$-\langle x_{h}-x^{*},\operatorname{\mathbb{E}}_{h}\left[g_{h}-P(x_{h}-x^{*})\right]\rangle$ $\leq\frac{\lambda_{P}}{4}\|x_{h}-x^{*}\|^{2}+\frac{1}{\lambda_{P}}\left\|\operatorname{\mathbb{E}}_{h}[g_{h}-P(x_{h}-x^{*})]\right\|^{2}$
$\leq\frac{\lambda_{P}}{4}\|x_{h}-x^{*}\|^{2}+\dfrac{1}{\lambda_{P}}\left\|\left\{\operatorname{\mathbb{E}}_{h}[\hat{P}_{h}]-P\right\}x_{h}+\bigg{\{}q-\operatorname{\mathbb{E}}_{h}\left[\hat{q}_{h}\right]\bigg{\}}\right\|^{2}$
$\leq\frac{\lambda_{P}}{4}\|x_{h}-x^{*}\|^{2}+\frac{2\delta_{P}^{2}\|x_{h}\|^{2}+2\delta_{q}^{2}}{\lambda_{P}}$
$\leq\frac{\lambda_{P}}{4}\|x_{h}-x^{*}\|^{2}+\frac{4\delta_{P}^{2}\|x_{h}-x^{*}\|^{2}+4\delta_{P}^{2}\lambda_{P}^{-2}\Lambda_{q}^{2}+2\delta_{q}^{2}}{\lambda_{P}}$ (40)
where the last inequality follows from $\left\lVert x^{*}\right\rVert^{2}=\left\lVert P^{-1}q\right\rVert^{2}\leq\lambda_{P}^{-2}\Lambda_{q}^{2}$. Substituting the above bounds in (B),
$\operatorname{\mathbb{E}}_{h}\left[\|x_{h+1}-x^{*}\|^{2}\right]$ $\leq\left(1-\frac{3\bar{\beta}\lambda_{P}}{2}+\dfrac{8\bar{\beta}\delta_{P}^{2}}{\lambda_{P}}+6\bar{\beta}^{2}\sigma_{P}^{2}+2\bar{\beta}^{2}\Lambda_{P}^{2}\right)\|x_{h}-x^{*}\|^{2}+\dfrac{4\bar{\beta}}{\lambda_{P}}\left[2\delta_{P}^{2}\lambda_{P}^{-2}\Lambda_{q}^{2}+\delta_{q}^{2}\right]$
$\hskip 28.45274pt+6\bar{\beta}^{2}\left[\lambda_{P}^{-2}\Lambda_{q}^{2}\sigma_{P}^{2}+\sigma_{q}^{2}\right]$
For $\delta_{P}\leq\lambda_{P}/8$, and $\bar{\beta}\leq\lambda_{P}/[4(6\sigma_{P}^{2}+2\Lambda_{P}^{2})]$, we can modify the above inequality to the following.
$\operatorname{\mathbb{E}}_{h}[\|x_{h+1}-x^{*}\|^{2}]\leq\left(1-{\beta\lambda_{P}}\right)\|x_{h}-x^{*}\|^{2}+\dfrac{4\bar{\beta}}{\lambda_{P}}\left[2\delta_{P}^{2}\lambda_{P}^{-2}\Lambda_{q}^{2}+\delta_{q}^{2}\right]+6\bar{\beta}^{2}\left[\lambda_{P}^{-2}\Lambda_{q}^{2}\sigma_{P}^{2}+\sigma_{q}^{2}\right]$
Taking expectation on both sides and unrolling the recursion yields
$\operatorname{\mathbb{E}}[\|x_{H}-x^{*}\|^{2}]$
$\leq\left(1-{\bar{\beta}\lambda_{P}}\right)^{H}\operatorname{\mathbb{E}}\|x_{0}-x^{*}\|^{2}+\sum_{h=0}^{H-1}\left(1-{\bar{\beta}\lambda_{P}}\right)^{h}\left\{\dfrac{4\bar{\beta}}{\lambda_{P}}\left[2\delta_{P}^{2}\lambda_{P}^{-2}\Lambda_{q}^{2}+\delta_{q}^{2}\right]+6\bar{\beta}^{2}\left[\lambda_{P}^{-2}\Lambda_{q}^{2}\sigma_{P}^{2}+\sigma_{q}^{2}\right]\right\}$
$\leq\exp\left(-{H\bar{\beta}\lambda_{P}}\right)\operatorname{\mathbb{E}}\|x_{0}-x^{*}\|^{2}+\frac{1}{\bar{\beta}\lambda_{P}}\left\{\dfrac{4\bar{\beta}}{\lambda_{P}}\left[2\delta_{P}^{2}\lambda_{P}^{-2}\Lambda_{q}^{2}+\delta_{q}^{2}\right]+6\bar{\beta}^{2}\left[\lambda_{P}^{-2}\Lambda_{q}^{2}\sigma_{P}^{2}+\sigma_{q}^{2}\right]\right\}$
$=\exp\left(-{H\bar{\beta}\lambda_{P}}\right)\operatorname{\mathbb{E}}\|x_{0}-x^{*}\|^{2}+\bigg{\{}4\lambda_{P}^{-2}\left[2\delta_{P}^{2}\lambda_{P}^{-2}\Lambda_{q}^{2}+\delta_{q}^{2}\right]+6\bar{\beta}\lambda_{P}^{-1}\left[\lambda_{P}^{-2}\Lambda_{q}^{2}\sigma_{P}^{2}+\sigma_{q}^{2}\right]\bigg{\}}$
Set $\bar{\beta}=\frac{2\log H}{\lambda_{P}H}$ to obtain the following result.
$\operatorname{\mathbb{E}}\left[\|x_{H}-x^{*}\|^{2}\right]\leq\frac{1}{H^{2}}\operatorname{\mathbb{E}}\left[\|x_{0}-x^{*}\|^{2}\right]+\operatorname*{\mathcal{O}}\left(\frac{\log H}{H}\underbrace{\bigg{\{}\lambda_{P}^{-4}\Lambda_{q}^{2}\sigma_{P}^{2}+\lambda_{P}^{-2}\sigma_{q}^{2}\bigg{\}}}_{R_{0}}+\underbrace{\lambda_{P}^{-2}\bigg{[}\delta_{P}^{2}\lambda_{P}^{-2}\Lambda_{q}^{2}+\delta_{q}^{2}\bigg{]}}_{R_{1}}\right)$ (41)
Note that, for consistency, we must have $\log H/H\leq\lambda_{P}^{2}/[8(6\sigma_{P}^{2}+2\Lambda_{P}^{2})]$. To prove the second statement, observe that we have the following recursion.
$\|\operatorname{\mathbb{E}}$ $[x_{h+1}]-x^{*}\|^{2}=\|\operatorname{\mathbb{E}}[x_{h}]-\bar{\beta}\operatorname{\mathbb{E}}[{g}_{h}]-x^{*}\|^{2}$
$=\|\operatorname{\mathbb{E}}[x_{h}]-x^{*}\|^{2}-2\bar{\beta}\langle\operatorname{\mathbb{E}}[x_{h}]-x^{*},\operatorname{\mathbb{E}}[{g}_{h}]\rangle+\bar{\beta}^{2}\|\operatorname{\mathbb{E}}[g_{h}]\|^{2}$
$\overset{}{=}\|\operatorname{\mathbb{E}}[x_{h}]-x^{*}\|^{2}-2\bar{\beta}\langle\operatorname{\mathbb{E}}[x_{h}]-x^{*},P(\operatorname{\mathbb{E}}[x_{h}]-x^{*})\rangle-2\bar{\beta}\langle\operatorname{\mathbb{E}}[x_{h}]-x^{*},\operatorname{\mathbb{E}}[{g}_{h}]-P(\operatorname{\mathbb{E}}[x_{h}]-x^{*})\rangle+\bar{\beta}^{2}\|\operatorname{\mathbb{E}}[g_{h}]\|^{2}$
$\overset{(a)}{\leq}\|\operatorname{\mathbb{E}}[x_{h}]-x^{*}\|^{2}-2\bar{\beta}\lambda_{P}\|\operatorname{\mathbb{E}}[x_{h}]-x^{*}\|^{2}-2\bar{\beta}\langle\operatorname{\mathbb{E}}[x_{h}]-x^{*},\operatorname{\mathbb{E}}[{g}_{h}]-P(\operatorname{\mathbb{E}}[x_{h}]-x^{*})\rangle$
$\hskip 28.45274pt+2\bar{\beta}^{2}\|\operatorname{\mathbb{E}}[g_{h}]-P(\operatorname{\mathbb{E}}[x_{h}]-x^{*})\|^{2}+2\bar{\beta}^{2}\|P(\operatorname{\mathbb{E}}[x_{h}]-x^{*})\|^{2}$
$\overset{(b)}{\leq}\|\operatorname{\mathbb{E}}[x_{h}]-x^{*}\|^{2}-2\bar{\beta}\lambda_{P}\|\operatorname{\mathbb{E}}[x_{h}]-x^{*}\|^{2}-2\bar{\beta}\langle\operatorname{\mathbb{E}}[x_{h}]-x^{*},\operatorname{\mathbb{E}}[g_{h}]-P(\operatorname{\mathbb{E}}[x_{h}]-x^{*})\rangle$
$\hskip 28.45274pt+2\bar{\beta}^{2}\|\operatorname{\mathbb{E}}[g_{h}]-P(\operatorname{\mathbb{E}}[x_{h}]-x^{*})\|^{2}+2\Lambda_{P}^{2}\bar{\beta}^{2}\|\operatorname{\mathbb{E}}[x_{h}]-x^{*}\|^{2}$
$\leq(1-2\bar{\beta}\lambda_{P}+2\Lambda_{P}^{2}\bar{\beta}^{2})\|\operatorname{\mathbb{E}}[x_{h}]-x^{*}\|^{2}-2\bar{\beta}\langle\operatorname{\mathbb{E}}[x_{h}]-x^{*},\operatorname{\mathbb{E}}[g_{h}]-P(\operatorname{\mathbb{E}}[x_{h}]-x^{*})\rangle$
$\hskip 28.45274pt+2\bar{\beta}^{2}\left\|\operatorname{\mathbb{E}}[g_{h}]-P(\operatorname{\mathbb{E}}[x_{h}]-x^{*})\right\|^{2}$ (42)
where $(a)$ and $(b)$ follow from $\lambda_{P}\leq\left\lVert P\right\rVert\leq\Lambda_{P}$. The third term in the last line of (B) can be bounded as follows.
$\|\operatorname{\mathbb{E}}[g_{h}]-P(\operatorname{\mathbb{E}}[x_{h}]-x^{*})\|^{2}$ $=\left\|\operatorname{\mathbb{E}}\left[(\hat{P}_{h}-P)(x_{h}-x^{*})\right]+(\operatorname{\mathbb{E}}[\hat{P}_{h}]-P)x^{*}+(q-\operatorname{\mathbb{E}}[\hat{q}_{h}])\right\|^{2}$
$\leq 3\operatorname{\mathbb{E}}\left[\|\operatorname{\mathbb{E}}_{h}[\hat{P}_{h}]-P\|^{2}\|x_{h}-x^{*}\|^{2}\right]+3\operatorname{\mathbb{E}}\left[\|\operatorname{\mathbb{E}}_{h}[\hat{P}_{h}]-P\|^{2}\right]\|x^{*}\|^{2}+3\left\|q-\operatorname{\mathbb{E}}[\hat{q}_{h}]\right\|^{2}$
$\leq 3\delta_{P}^{2}\operatorname{\mathbb{E}}\left[\|x_{h}-x^{*}\|^{2}\right]+3\lambda_{P}^{-2}\Lambda_{q}^{2}\delta_{P}^{2}+3\bar{\delta}_{q}^{2}$
$\overset{(a)}{\leq}3\delta_{P}^{2}\bigg{\{}\operatorname{\mathbb{E}}\left[\|x_{0}-x^{*}\|^{2}\right]+\mathcal{O}\left(R_{0}+R_{1}\right)\bigg{\}}+3\lambda_{P}^{-2}\Lambda_{q}^{2}\delta_{P}^{2}+3\bar{\delta}_{q}^{2}$
where $(a)$ follows from (41). The second term in the last line of (B) can be bounded as follows.
$-\langle\operatorname{\mathbb{E}}[x_{h}]$ $-x^{*},\operatorname{\mathbb{E}}_{h}\left[\operatorname{\mathbb{E}}[g_{h}]-P(\operatorname{\mathbb{E}}[x_{h}]-x^{*})\right]\rangle$
$\leq\frac{\lambda_{P}}{4}\|\operatorname{\mathbb{E}}[x_{h}]-x^{*}\|^{2}+\frac{1}{\lambda_{P}}\left\|\operatorname{\mathbb{E}}[g_{h}]-P(\operatorname{\mathbb{E}}[x_{h}]-x^{*})\right\|^{2}$
$\leq\frac{\lambda_{P}}{4}\|\operatorname{\mathbb{E}}[x_{h}]-x^{*}\|^{2}+\dfrac{3}{\lambda_{P}}\left[\delta_{P}^{2}\bigg{\{}\operatorname{\mathbb{E}}\left[\|x_{0}-x^{*}\|^{2}\right]+\mathcal{O}\left(R_{0}+R_{1}\right)\bigg{\}}+\lambda_{P}^{-2}\Lambda_{q}^{2}\delta_{P}^{2}+\bar{\delta}_{q}^{2}\right]$
Substituting the above bounds in (B), we obtain the following recursion.
$\|\operatorname{\mathbb{E}}[x_{h+1}]-x^{*}\|^{2}$ $\leq\left(1-\dfrac{3\bar{\beta}\lambda_{P}}{2}+2\Lambda_{P}^{2}\bar{\beta}^{2}\right)\|\operatorname{\mathbb{E}}[x_{h}]-x^{*}\|^{2}$
$\hskip 28.45274pt+6\bar{\beta}\left(\bar{\beta}+\dfrac{1}{\lambda_{P}}\right)\left[\delta_{P}^{2}\bigg{\{}\operatorname{\mathbb{E}}\left[\|x_{0}-x^{*}\|^{2}\right]+\mathcal{O}\left(R_{0}+R_{1}\right)\bigg{\}}+\lambda_{P}^{-2}\Lambda_{q}^{2}\delta_{P}^{2}+\bar{\delta}_{q}^{2}\right]$
If $\beta<\lambda_{P}/(4\Lambda_{P}^{2})$, the above bound implies the following.
$\|\operatorname{\mathbb{E}}[x_{h+1}]-x^{*}\|^{2}$ $\leq\left(1-{\bar{\beta}\lambda_{P}}\right)\|\operatorname{\mathbb{E}}[x_{h}]-x^{*}\|^{2}$
$\hskip 28.45274pt+6\bar{\beta}\left(\bar{\beta}+\dfrac{1}{\lambda_{P}}\right)\left[\delta_{P}^{2}\bigg{\{}\operatorname{\mathbb{E}}\left[\|x_{0}-x^{*}\|^{2}\right]+\mathcal{O}\left(R_{0}+R_{1}\right)\bigg{\}}+\lambda_{P}^{-2}\Lambda_{q}^{2}\delta_{P}^{2}+\bar{\delta}_{q}^{2}\right]$
Unrolling the above recursion, we obtain
$\|\operatorname{\mathbb{E}}[x_{H}]-x^{*}\|^{2}\leq\left(1-{\bar{\beta}\lambda_{P}}\right)^{H}\|\operatorname{\mathbb{E}}[x_{0}]-x^{*}\|^{2}$
$\hskip 42.67912pt+\sum_{h=0}^{H-1}6\left(1-{\bar{\beta}\lambda_{P}}\right)^{h}\bar{\beta}\left(\bar{\beta}+\dfrac{1}{\lambda_{P}}\right)\left[\delta_{P}^{2}\bigg{\{}\operatorname{\mathbb{E}}\left[\|x_{0}-x^{*}\|^{2}\right]+\mathcal{O}\left(R_{0}+R_{1}\right)\bigg{\}}+\lambda_{P}^{-2}\Lambda_{q}^{2}\delta_{P}^{2}+\bar{\delta}_{q}^{2}\right]$
$\leq\exp\left(-{H\bar{\beta}\lambda_{P}}\right)\|\operatorname{\mathbb{E}}[x_{0}]-x^{*}\|^{2}+\dfrac{6}{\lambda_{P}}\left(\bar{\beta}+\dfrac{1}{\lambda_{P}}\right)\left[\delta_{P}^{2}\bigg{\{}\operatorname{\mathbb{E}}\left[\|x_{0}-x^{*}\|^{2}\right]+\mathcal{O}\left(R_{0}+R_{1}\right)\bigg{\}}+\lambda_{P}^{-2}\Lambda_{q}^{2}\delta_{P}^{2}+\bar{\delta}_{q}^{2}\right]$
Substituting $\bar{\beta}=2\log H/(\lambda_{P}H)$, we finally arrive at the following result.
$\|\operatorname{\mathbb{E}}[x_{H}]-x^{*}\|^{2}\leq\dfrac{1}{H^{2}}\|\operatorname{\mathbb{E}}[x_{0}]-x^{*}\|^{2}+6\left(1+\dfrac{2\log H}{H}\right)\left[\lambda_{P}^{-2}\delta_{P}^{2}\bigg{\{}\operatorname{\mathbb{E}}\left[\|x_{0}-x^{*}\|^{2}\right]+\mathcal{O}\left(R_{0}+R_{1}\right)\bigg{\}}+\bar{R}_{1}\right]$
where $\bar{R}_{1}=\lambda_{P}^{-2}\big{[}\delta_{P}^{2}\lambda_{P}^{-2}\Lambda_{q}^{2}+\bar{\delta}_{q}^{2}\big{]}$. This concludes the result.
Appendix C Properties of the MLMC Estimates
This section provides some guarantees on the error of the MLMC estimator. This is similar to the results available in [Dorfman and Levy, 2022, Suttle et al., 2023, Beznosikov et al., 2023], although [Dorfman and Levy, 2022, Beznosikov et al., 2023] consider the case of unbiased estimates while our results deal with biased estimates.
Lemma 5.
Consider a time-homogeneous, ergodic Markov chain $(Z_{t})_{t\geq 0}$ with a unique stationary distribution $d_{Z}$ and a mixing time $t_{\mathrm{mix}}$. Assume that $\nabla F(x)$ is an arbitrary gradient and $\nabla F(x,Z)$ denotes an estimate of $\nabla F(x)$. Let $\|\operatorname{\mathbb{E}}_{d_{Z}}[\nabla F(x,Z)]-\nabla F(x)\|^{2}\leq\delta^{2}$ and $\|\nabla F(x,Z_{t})-\operatorname{\mathbb{E}}_{d_{Z}}\left[\nabla F(x,Z)\right]\|^{2}\leq\sigma^{2}$ for all $t\geq 0$. If $Q\sim\mathrm{Geom}(1/2)$, then the following MLMC estimator
$g_{\mathrm{MLMC}}=g^{0}+\begin{cases}\textstyle{2^{Q}\left(g^{Q}-g^{Q-1}\right)},&\text{if}~{}2^{Q}\leq T_{\max}\\ 0,&\text{otherwise}\end{cases}\text{~{}~{}where}~{}g^{j}=2^{-j}\sum\nolimits_{t=1}^{2^{j}}\nabla F(x,Z_{t})$ (43)
satisfies the inequalities stated below.
  1. (a)
    $\operatorname{\mathbb{E}}[g_{\mathrm{MLMC}}]=\operatorname{\mathbb{E}}[g^{\lfloor\log T_{\max}\rfloor}]$
  2. (b)
    $\operatorname{\mathbb{E}}[\|\nabla F(x)-g_{\mathrm{MLMC}}\|^{2}]\leq\mathcal{O}\left(\sigma^{2}t_{\mathrm{mix}}\log_{2}T_{\max}+\delta^{2}\right)$
  3. (c)
    $\|\nabla F(x)-\operatorname{\mathbb{E}}[g_{\mathrm{MLMC}}]\|^{2}\leq\operatorname*{\mathcal{O}}\left(\sigma^{2}t_{\mathrm{mix}}T_{\max}^{-1}+\delta^{2}\right)$
Before proceeding to the proof, we state a useful lemma.
Lemma 6 (Lemma 1, [Beznosikov et al., 2023]).
Consider the same setup as in Lemma 5. The following inequality holds.
${\operatorname{\mathbb{E}}\left[\left\lVert\dfrac{1}{N}\sum\limits_{i=1}^{N}\nabla F(x,Z_{i})-\operatorname{\mathbb{E}}_{d_{Z}}\left[\nabla F(x,Z)\right]\right\rVert^{2}\right]\leq\dfrac{C_{1}t_{\mathrm{mix}}}{N}\sigma^{2}}\,$ (44)
where $N$ is a constant, $C_{1}=16(1+\frac{1}{\ln^{2}{4}})$, and the expectation is over the distribution of $\{Z_{i}\}_{i=1}^{N}$ emanating from any arbitrary initial distribution.
Proof of Lemma 5.
The statement $(a)$ can be proven as follows.
$\operatorname{\mathbb{E}}[g_{\mathrm{MLMC}}]$ $=\operatorname{\mathbb{E}}[g^{0}]+\sum\limits_{j=1}^{\lfloor\log_{2}T_{\max}\rfloor}\Pr\{Q=j\}\cdot 2^{j}\operatorname{\mathbb{E}}[g^{j}-g^{j-1}]$
$=\operatorname{\mathbb{E}}[g^{0}]+\sum\limits_{j=1}^{\lfloor\log_{2}T_{\max}\rfloor}\operatorname{\mathbb{E}}[g^{j}-g^{j-1}]=\operatorname{\mathbb{E}}[g^{\lfloor\log_{2}T_{\max}\rfloor}]\,$
For the proof of $(b)$, notice that
$\operatorname{\mathbb{E}}\left[\left\|\operatorname{\mathbb{E}}_{d_{Z}}\left[\nabla F(x,Z)\right]-g_{\mathrm{MLMC}}\right\|^{2}\right]$
$\leq 2\operatorname{\mathbb{E}}\left[\left\|\operatorname{\mathbb{E}}_{d_{Z}}\left[\nabla F(x_{t})\right]-g^{0}\right\|^{2}\right]+2\operatorname{\mathbb{E}}\left[\left\|g_{\mathrm{MLMC}}-g^{0}\right\|^{2}\right]$
$=2\operatorname{\mathbb{E}}\left[\left\|\operatorname{\mathbb{E}}_{d_{Z}}\left[\nabla F(x_{t})\right]-g^{0}\right\|^{2}\right]+2\sum\nolimits_{j=1}^{\lfloor\log_{2}T_{\max}\rfloor}\Pr\{Q=j\}\cdot 4^{j}\operatorname{\mathbb{E}}\left[\left\|g^{j}-g^{j-1}\right\|^{2}\right]$
$=2\operatorname{\mathbb{E}}\left[\left\|\operatorname{\mathbb{E}}_{d_{Z}}\left[\nabla F(x_{t})\right]-g^{0}\right\|^{2}\right]+2\sum\nolimits_{j=1}^{\lfloor\log_{2}T_{\max}\rfloor}2^{j}\operatorname{\mathbb{E}}\left[\left\|g^{j}-g^{j-1}\right\|^{2}\right]$
$\leq 2\operatorname{\mathbb{E}}\left[\left\|\operatorname{\mathbb{E}}_{d_{Z}}\left[\nabla F(x_{t})\right]-g^{0}\right\|^{2}\right]$
$\hskip 28.45274pt+4\sum\nolimits_{j=1}^{\lfloor\log_{2}T_{\max}\rfloor}2^{j}\left(\operatorname{\mathbb{E}}\left[\left\|\operatorname{\mathbb{E}}_{d_{z}}\left[\nabla F(x,Z)\right]-g^{j-1}\right\|^{2}\right]+\operatorname{\mathbb{E}}\left[\left\|g^{j}-\operatorname{\mathbb{E}}_{d_{Z}}\left[\nabla F(x,Z)\right]\right\|^{2}\right]\right)\,$
$\overset{(a)}{\leq}C_{1}t_{\mathrm{mix}}\sigma^{2}\left[2+4\sum\nolimits_{j=1}^{\lfloor\log_{2}T_{\max}\rfloor}2^{j}\left(\dfrac{1}{2^{j-1}}+\dfrac{1}{2^{j}}\right)\right]$
$=\mathcal{O}\left(\sigma^{2}t_{\mathrm{mix}}\log_{2}T_{\max}\right)$
where $(a)$ follows from Lemma 6. Using this result, we obtain the following.
$\operatorname{\mathbb{E}}\left[\left\|\nabla F(x)-g_{\mathrm{MLMC}}\right\|^{2}\right]$ $\leq 2\operatorname{\mathbb{E}}\left[\left\|\nabla F(x)-\operatorname{\mathbb{E}}_{d_{z}}\left[\nabla F(x,Z)\right]\right\|^{2}\right]+2\operatorname{\mathbb{E}}\left[\left\|\operatorname{\mathbb{E}}_{d_{z}}\left[\nabla F(x,Z)\right]-g_{\mathrm{MLMC}}\right\|^{2}\right]$
$\leq\mathcal{O}\left(\sigma^{2}t_{\mathrm{mix}}\log_{2}T_{\max}+\delta^{2}\right)$
This completes the proof of statement $(b)$. For part $(c)$, we have
$\begin{split}\left\|\nabla F(x)-\operatorname{\mathbb{E}}\left[g_{\mathrm{MLMC}}\right]\right\|^{2}&\leq 2\left\|\nabla F(x)-\operatorname{\mathbb{E}}_{d_{Z}}\left[\nabla F(x,Z)\right]\right\|^{2}+2\left\|\operatorname{\mathbb{E}}_{d_{Z}}\left[\nabla F(x,Z)\right]-\operatorname{\mathbb{E}}\left[g_{\mathrm{MLMC}}\right]\right\|^{2}\\ &\leq 2\delta^{2}+2\left\|\operatorname{\mathbb{E}}_{d_{Z}}\left[\nabla F(x,Z)\right]-\operatorname{\mathbb{E}}[g^{\lfloor\log_{2}T_{\max}\rfloor}]\right\|^{2}\overset{(a)}{\leq}2\delta^{2}+\frac{2C_{1}t_{\mathrm{mix}}}{T_{\max}}\sigma^{2}\end{split}$ (45)
where $(a)$ follows from Lemma 6. This concludes the proof of Lemma 5. ∎
Appendix D Proof of Lemma 2
Fix an outer loop instant $k$ and an inner loop instant $h\in\{H,\cdots,2H-1\}$. Recall the definition of $\hat{A}_{\mathbf{u}}(\theta_{k},\cdot)$ from (19). The following inequalities hold for any $\theta_{k}$ and $z_{t}^{kh}\in\mathcal{S}\times\mathcal{A}\times\mathcal{S}$.
$\operatorname{\mathbb{E}}_{\theta_{k}}\left[\hat{A}_{\mathbf{u}}(\theta_{k},z)\right]\overset{(a)}{=}F(\theta_{k}),~{}\text{and}~{}\left\lVert\hat{A}_{\mathbf{u}}(\theta_{k},z_{t}^{kh})-\operatorname{\mathbb{E}}_{\theta_{k}}\left[\hat{A}_{\mathbf{u}}(\theta_{k},z)\right]\right\rVert^{2}\overset{(b)}{\leq}2G_{1}^{4}$
where $\operatorname{\mathbb{E}}_{\theta_{k}}$ denotes the expectation over the distribution of $z=(s,a,s^{\prime})$ where $(s,a)\sim\nu^{\pi_{\theta_{k}}}$, $s^{\prime}\sim P(\cdot|s,a)$. The equality $(a)$ follows from the definition of the Fisher matrix, and $(b)$ is a consequence of Assumption 5. Statements (a) and (b), therefore, directly follow from Lemma 5.
To prove the other statements, recall the definition of $\hat{b}_{\mathbf{u}}(\theta_{k},\xi_{k},\cdot)$ from (19). Observe the following relations for arbitrary $\theta_{k},\xi_{k}$.
$\operatorname{\mathbb{E}}_{\theta_{k}}\left[\hat{b}_{\mathbf{u}}(\theta_{k},\xi_{k},z)\right]$ $-\nabla_{\theta}J(\theta_{k})=\operatorname{\mathbb{E}}_{\theta_{k}}\left[\bigg{\{}r(s,a)-\eta_{k}+\langle\phi(s^{\prime})-\phi(s),\zeta_{k}\rangle\bigg{\}}\nabla_{\theta_{k}}\log_{\pi_{\theta_{k}}}(a|s)\right]-\nabla_{\theta}J(\theta_{k})$
=$\overset{(a)}{=}\underbrace{\operatorname{\mathbb{E}}_{\theta_{k}}\left[\bigg{\{}\eta_{k}^{*}-\eta_{k}+\langle\phi(s^{\prime})-\phi(s),\zeta_{k}-\zeta_{k}^{*}\rangle\bigg{\}}\nabla_{\theta_{k}}\log_{\pi_{\theta_{k}}}(a|s)\right]}_{T_{0}}+$
$\hskip 28.45274pt+\underbrace{\operatorname{\mathbb{E}}_{\theta_{k}}\left[\bigg{\{}V^{\pi_{\theta_{k}}}(s)-\langle\phi(s),\zeta_{k}^{*}\rangle+\langle\phi(s^{\prime}),\zeta_{k}^{*}\rangle-V^{\pi_{\theta_{k}}}(s^{\prime})\bigg{\}}\nabla_{\theta_{k}}\log_{\pi_{\theta_{k}}}(a|s)\right]}_{T_{1}}$
$\hskip 28.45274pt+\underbrace{\operatorname{\mathbb{E}}_{\theta_{k}}\left[\bigg{\{}V^{\pi_{\theta_{k}}}(s^{\prime})-\eta_{k}^{*}+r(s,a)-V^{\pi_{\theta_{k}}}(s)\bigg{\}}\nabla_{\theta_{k}}\log_{\pi_{\theta_{k}}}(a|s)\right]-\nabla_{\theta}J(\theta_{k})}_{T_{2}}$
In $(a)$, we have used the notation that $\xi_{k}^{*}=[\eta_{k}^{*},\zeta_{k}^{*}]^{\top}$. Observe that
$\left\lVert T_{0}\right\rVert^{2}$ =$\overset{(a)}{=}\operatorname*{\mathcal{O}}\left(G_{1}^{2}\left\lVert\xi_{k}-\xi_{k}^{*}\right\rVert^{2}\right),~{}\left\lVert T_{1}\right\rVert^{2}\overset{(b)}{=}\operatorname*{\mathcal{O}}\left(G_{1}^{2}\epsilon_{\mathrm{app}}\right),~{}\text{and}~{}T_{2}\overset{(c)}{=}0$ (46)
where $(a)$ follows from Assumption 5 and the boundedness of the feature map, $\phi$ while $(b)$ is a consequence of Assumption 5 and 2. Finally, $(c)$ is an application of Bellman’s equation. We get,
$\left\lVert\operatorname{\mathbb{E}}_{\theta_{k}}\left[\hat{b}_{\mathbf{u}}(\theta_{k},\xi_{k},z)\right]-\nabla_{\theta}J(\theta_{k})\right\rVert^{2}\leq\delta_{\mathbf{u},k}^{2}=\operatorname*{\mathcal{O}}\left(G_{1}^{2}\left\lVert\xi_{k}-\xi_{k}^{*}\right\rVert^{2}+G_{1}^{2}\epsilon_{\mathrm{app}}\right)$ (47)
Moreover, observe that, for arbitrary $z_{t}^{kh}\in\mathcal{S}\times\mathcal{A}\times\mathcal{S}$
$\left\lVert\hat{b}_{\mathbf{u}}(\theta_{k},\xi_{k},z_{t}^{kh})-\operatorname{\mathbb{E}}_{\theta_{k}}\left[\hat{b}_{\mathbf{u}}(\theta_{k},\xi_{k},z_{t})\right]\right\rVert^{2}\overset{(a)}{\leq}\sigma_{\mathbf{u},k}^{2}=\operatorname*{\mathcal{O}}\left(G_{1}^{2}\left\lVert\xi_{k}\right\rVert^{2}\right)$ (48)
where $(a)$ follows from Assumption 5 and the boundedness of the feature map, $\phi$. We can, therefore, conclude statements (c) and (d) by applying (47) and (48) in Lemma 5. To prove the statement (e), note that
$\operatorname{\mathbb{E}}_{\theta_{k}}\left[\operatorname{\mathbb{E}}_{k}\left[\hat{b}_{\mathbf{u}}(\theta_{k},\xi_{k},z)\right]\right]$ $-\nabla_{\theta}J(\theta_{k})=\operatorname{\mathbb{E}}_{\theta_{k}}\left[\bigg{\{}r(s,a)-\operatorname{\mathbb{E}}_{k}[\eta_{k}]+\langle\phi(s^{\prime})-\phi(s),\operatorname{\mathbb{E}}_{k}[\zeta_{k}]\rangle\bigg{\}}\nabla_{\theta_{k}}\log_{\pi_{\theta_{k}}}(a|s)\right]-\nabla_{\theta}J(\theta_{k})$
=$\overset{(a)}{=}\underbrace{\operatorname{\mathbb{E}}_{\theta_{k}}\left[\bigg{\{}\eta_{k}^{*}-\operatorname{\mathbb{E}}_{k}[\eta_{k}]+\langle\phi(s^{\prime})-\phi(s),\operatorname{\mathbb{E}}_{k}[\zeta_{k}]-\zeta_{k}^{*}\rangle\bigg{\}}\nabla_{\theta_{k}}\log_{\pi_{\theta_{k}}}(a|s)\right]}_{T_{0}}+$
$\hskip 28.45274pt+\underbrace{\operatorname{\mathbb{E}}_{\theta_{k}}\left[\bigg{\{}V^{\pi_{\theta_{k}}}(s)-\langle\phi(s),\zeta_{k}^{*}\rangle+\langle\phi(s^{\prime}),\zeta_{k}^{*}\rangle-V^{\pi_{\theta_{k}}}(s^{\prime})\bigg{\}}\nabla_{\theta_{k}}\log_{\pi_{\theta_{k}}}(a|s)\right]}_{T_{1}}$
$\hskip 28.45274pt+\underbrace{\operatorname{\mathbb{E}}_{\theta_{k}}\left[\bigg{\{}V^{\pi_{\theta_{k}}}(s^{\prime})-\eta_{k}^{*}+r(s,a)-V^{\pi_{\theta_{k}}}(s)\bigg{\}}\nabla_{\theta_{k}}\log_{\pi_{\theta_{k}}}(a|s)\right]-\nabla_{\theta}J(\theta_{k})}_{T_{2}}$
Observe the following bounds.
$\left\lVert T_{0}\right\rVert^{2}$ =$\overset{(a)}{=}\operatorname*{\mathcal{O}}\left(G_{1}^{2}\left\lVert\operatorname{\mathbb{E}}_{k}[\xi_{k}]-\xi_{k}^{*}\right\rVert^{2}\right),~{}\left\lVert T_{1}\right\rVert^{2}\overset{(b)}{=}\operatorname*{\mathcal{O}}\left(G_{1}^{2}\epsilon_{\mathrm{app}}\right),~{}\text{and}~{}T_{2}\overset{(c)}{=}0$ (49)
where $(a)$ follows from Assumption 5 and the boundedness of the feature map, $\phi$ while $(b)$ is a consequence of Assumption 5 and 2. Finally, $(c)$ is an application of Bellman’s equation. We get,
$\left\lVert\operatorname{\mathbb{E}}_{\theta_{k}}\left[\operatorname{\mathbb{E}}_{k}\left[\hat{b}_{\mathbf{u}}(\theta_{k},\xi_{k},z)\right]\right]-\nabla_{\theta}J(\theta_{k})\right\rVert^{2}\leq\bar{\delta}_{\mathbf{u},k}^{2}=\operatorname*{\mathcal{O}}\left(G_{1}^{2}\left\lVert\operatorname{\mathbb{E}}_{k}[\xi_{k}]-\xi_{k}^{*}\right\rVert^{2}+G_{1}^{2}\epsilon_{\mathrm{app}}\right)$ (50)
Using the above bound, we deduce the following.
$\left\lVert\operatorname{\mathbb{E}}_{k}\left[\hat{b}^{\mathrm{MLMC}}_{\mathbf{u},k,h}(\theta_{k},\xi_{k})\right]-\nabla_{\theta}J(\theta_{k})\right\rVert^{2}$
$\leq 2\left\lVert\operatorname{\mathbb{E}}_{k}\left[\hat{b}^{\mathrm{MLMC}}_{\mathbf{u},k,h}(\theta_{k},\xi_{k})\right]-\operatorname{\mathbb{E}}_{\theta_{k}}\left[\operatorname{\mathbb{E}}_{k}\left[\hat{b}_{\mathbf{u}}(\theta_{k},\xi_{k},z)\right]\right]\right\rVert^{2}+2\left\lVert\operatorname{\mathbb{E}}_{\theta_{k}}\left[\operatorname{\mathbb{E}}_{k}\left[\hat{b}_{\mathbf{u}}(\theta_{k},\xi_{k},z)\right]\right]-\nabla_{\theta}J(\theta_{k})\right\rVert^{2}$
$\overset{(a)}{\leq}2\operatorname{\mathbb{E}}_{k}\left\lVert\operatorname{\mathbb{E}}_{k,h}\left[\hat{b}^{\mathrm{MLMC}}_{\mathbf{u},k,h}(\theta_{k},\xi_{k})\right]-\operatorname{\mathbb{E}}_{\theta_{k}}\left[\hat{b}_{\mathbf{u}}(\theta_{k},\xi_{k},z)\right]\right\rVert^{2}+\operatorname*{\mathcal{O}}\left(\bar{\delta}_{\mathbf{u},k}^{2}\right)\overset{(b)}{\leq}\operatorname*{\mathcal{O}}\left(t_{\mathrm{mix}}T_{\max}^{-1}\bar{\sigma}_{\mathbf{u},k}^{2}+\bar{\delta}_{\mathbf{u},k}^{2}\right)$
where $(a)$ follows from (50). Moreover, $(b)$ follows from Lemma 5(a), 6, and the definition of $\bar{\sigma}^{2}_{\mathbf{u},k}$. This concludes the proof of Lemma 2.
Appendix E Proof of Lemma 3
Recall the definitions of $\hat{A}_{\mathbf{v}}(\cdot)$ and $\hat{b}_{\mathbf{v}}(\cdot)$ given in (22) and (23) respectively. Note that the following equalities hold for any $\theta_{k}$.
$\operatorname{\mathbb{E}}_{\theta_{k}}\left[\hat{A}_{\mathbf{v}}(z)\right]=A_{\mathbf{v}}(\theta_{k}),~{}\text{and}~{}\operatorname{\mathbb{E}}_{\theta_{k}}\left[\hat{b}_{\mathbf{v}}(z)\right]=b_{\mathbf{v}}(\theta_{k})$ (51)
where $\operatorname{\mathbb{E}}_{\theta_{k}}$ denotes the expectation over the distribution of $z=(s,a,s^{\prime})$ where $(s,a)\sim\nu^{\pi_{\theta_{k}}}$, $s^{\prime}\sim P(\cdot|s,a)$. Also, for any $z=(s,a,s^{\prime})\in\mathcal{S}\times\mathcal{A}\times\mathcal{S}$, we have the following.
$\left\lVert\hat{A}_{\mathbf{v}}(z)\right\rVert\leq|c_{\beta}|+\left\lVert\phi(s)\right\rVert+\left\lVert\phi(s)(\phi(s)-\phi(s^{\prime}))^{\top}\right\rVert\overset{(a)}{\leq}c_{\beta}+3=\mathcal{O}(c_{\beta}),$ (52)
$\left\lVert\hat{b}_{\mathbf{v}}(z)\right\rVert\leq|c_{\beta}r(s,a)|+\left\lVert r(s,a)\phi(s)\right\rVert\overset{(b)}{\leq}c_{\beta}+1=\mathcal{O}(c_{\beta})$ (53)
where $(a)$, $(b)$ hold since $|r(s,a)|\leq 1$ and $\left\lVert{\phi(s)}\right\rVert\leq 1$, $\forall(s,a)\in\mathcal{S}\times\mathcal{A}$. Hence, for any $z_{t}^{kh}\in\mathcal{S}\times\mathcal{A}\times\mathcal{S}$, we have
$\left\lVert\hat{A}_{\mathbf{v}}(z_{t}^{kh})-\operatorname{\mathbb{E}}_{\theta_{k}}\left[\hat{A}_{\mathbf{v}}(z)\right]\right\rVert^{2}\leq\mathcal{O}(c_{\beta}^{2}),~{}\text{and}~{}\left\lVert\hat{b}_{\mathbf{v}}(z_{t}^{kh})-\operatorname{\mathbb{E}}_{\theta_{k}}\left[\hat{b}_{\mathbf{v}}(z)\right]\right\rVert^{2}\leq\mathcal{O}(c_{\beta}^{2})$
Combining the above results with Lemma 5 establishes the result.
Appendix F Proof of Theorem 3
We first state an important result regarding ergodic MPDs.
Lemma 7.
[Wei et al., 2020, Lemma 14] For any ergodic MDP with mixing time $t_{\mathrm{mix}}$, the following holds for any policy $\pi$.
$|A^{\pi}(s,a)|=\operatorname*{\mathcal{O}}\left(t_{\mathrm{mix}}\right),~{}\forall(s,a)$
If follows from Assumptions 5, 6, and Lemma 7 that
$\mu\leq\left\lVert F(\theta)\right\rVert\leq G_{1}^{2},~{}\text{and}~{}\left\lVert\nabla_{\theta}J(\theta)\right\rVert\leq\operatorname*{\mathcal{O}}\left(G_{1}t_{\mathrm{mix}}\right)$ (54)
where $\theta$ is any arbitrary policy parameter. Combining the above results with Lemma 2 and invoking Theorem 2, we arrive at the following.
$\operatorname{\mathbb{E}}_{k}\left[\left\lVert\omega_{k}-\omega_{k}^{*}\right\rVert^{2}\right]\leq\dfrac{1}{H^{2}}\left\lVert\omega_{H}^{k}-\omega_{k}^{*}\right\rVert^{2}+\tilde{\operatorname*{\mathcal{O}}}\left(\dfrac{R_{0}}{H}+R_{1}\right),$
$\left\lVert\operatorname{\mathbb{E}}_{k}[\omega_{k}]-\omega_{k}^{*}\right\rVert^{2}\leq\dfrac{1}{H^{2}}\left\lVert\omega^{k}_{H}-\omega_{k}^{*}\right\rVert^{2}+\operatorname*{\mathcal{O}}(\bar{R}_{1})+\operatorname*{\mathcal{O}}\left(\dfrac{G_{1}^{4}t_{\mathrm{mix}}}{\mu^{2}H^{2}}\left\{\left\lVert\omega_{H}^{k}-\omega_{k}^{*}\right\rVert^{2}+\tilde{\operatorname*{\mathcal{O}}}\left(R_{0}+R_{1}\right)\right\}\right)$
where the terms $R_{0},R_{1},\bar{R}_{1}$ are defined as follows.
$R_{0}$ $=\tilde{\operatorname*{\mathcal{O}}}\left(\mu^{-4}G_{1}^{6}t_{\mathrm{mix}}^{3}+\mu^{-2}G_{1}^{2}t_{\mathrm{mix}}\operatorname{\mathbb{E}}_{k}\left[\left\lVert\xi_{k}\right\rVert^{2}\right]+\mu^{-2}G_{1}^{2}\operatorname{\mathbb{E}}_{k}\left[\left\lVert\xi_{k}-\xi_{k}^{*}\right\rVert^{2}\right]+\mu^{-2}G_{1}^{2}\epsilon_{\mathrm{app}}\right),$
$R_{1}$ $=\operatorname*{\mathcal{O}}\left(H^{-2}\mu^{-4}G_{1}^{6}t^{3}_{\mathrm{mix}}+H^{-2}\mu^{-2}G_{1}^{2}t_{\mathrm{mix}}\operatorname{\mathbb{E}}_{k}\left[\left\lVert\xi_{k}\right\rVert^{2}\right]+\mu^{-2}G_{1}^{2}\operatorname{\mathbb{E}}_{k}\left[\left\lVert\xi_{k}-\xi_{k}^{*}\right\rVert^{2}\right]+\mu^{-2}G_{1}^{2}\epsilon_{\mathrm{app}}\right)$
$\bar{R}_{1}$ $=\operatorname*{\mathcal{O}}\left(H^{-2}\mu^{-4}G_{1}^{6}t^{3}_{\mathrm{mix}}+H^{-2}\mu^{-2}G_{1}^{2}t_{\mathrm{mix}}\operatorname{\mathbb{E}}_{k}\left[\left\lVert\xi_{k}\right\rVert^{2}\right]+\mu^{-2}G_{1}^{2}\left\lVert\operatorname{\mathbb{E}}_{k}[\xi_{k}]-\xi_{k}^{*}\right\rVert^{2}+\mu^{-2}G_{1}^{2}\epsilon_{\mathrm{app}}\right)$
Moreover, note that
$\operatorname{\mathbb{E}}_{k}\left[\left\lVert\xi_{k}\right\rVert^{2}\right]\leq 2\operatorname{\mathbb{E}}_{k}\left[\left\lVert\xi_{k}-\xi_{k}^{*}\right\rVert^{2}\right]+2\operatorname{\mathbb{E}}_{k}\left[\left\lVert\xi_{k}^{*}\right\rVert^{2}\right]\overset{(a)}{\leq}\operatorname*{\mathcal{O}}\left(\operatorname{\mathbb{E}}_{k}\left[\left\lVert\xi_{k}-\xi_{k}^{*}\right\rVert^{2}\right]+\lambda^{-2}c_{\beta}^{2}\right)$
where $(a)$ follows from (56) for sufficiently large $c_{\beta}$ and the definition that $\xi_{k}^{*}=[A_{\mathbf{v}}(\theta_{k})]^{-1}b_{\mathbf{v}}(\theta_{k})$. Hence,
$\operatorname{\mathbb{E}}_{k}\left[\left\lVert\omega_{k}-\omega_{k}^{*}\right\rVert^{2}\right]\leq\dfrac{1}{H^{2}}\left\lVert\omega_{H}^{k}-\omega_{k}^{*}\right\rVert^{2}+\tilde{\operatorname*{\mathcal{O}}}\left(\dfrac{1}{H}\bigg{\{}\mu^{-4}G_{1}^{6}t_{\mathrm{mix}}^{3}+\mu^{-2}\lambda^{-2}G_{1}^{2}c_{\beta}^{2}t_{\mathrm{mix}}\bigg{\}}\right)$
$\hskip 34.14322pt+\tilde{\operatorname*{\mathcal{O}}}\left(\mu^{-2}G_{1}^{2}\operatorname{\mathbb{E}}_{k}\left[\left\lVert\xi_{k}-\xi_{k}^{*}\right\rVert^{2}\right]+\mu^{-2}G_{1}^{2}\epsilon_{\mathrm{app}}\right),$
$\left\lVert\operatorname{\mathbb{E}}_{k}[\omega_{k}]-\omega_{k}^{*}\right\rVert^{2}\leq\operatorname*{\mathcal{O}}\left(\mu^{-2}G_{1}^{2}\left\lVert\operatorname{\mathbb{E}}_{k}[\xi_{k}]-\xi_{k}^{*}\right\rVert^{2}+\mu^{-2}G_{1}^{2}\epsilon_{\mathrm{app}}\right)$
$\hskip 34.14322pt+\tilde{\operatorname*{\mathcal{O}}}\left(\dfrac{G_{1}^{4}t_{\mathrm{mix}}}{\mu^{2}H^{2}}\left\{\left\lVert\omega_{H}^{k}-\omega_{k}^{*}\right\rVert^{2}+\mu^{-2}G_{1}^{2}t_{\mathrm{mix}}\operatorname{\mathbb{E}}_{k}\left[\left\lVert\xi_{k}-\xi_{k}^{*}\right\rVert^{2}\right]+\mu^{-4}G_{1}^{6}t_{\mathrm{mix}}^{3}+\mu^{-2}\lambda^{-2}G_{1}^{2}c_{\beta}^{2}t_{\mathrm{mix}}\right\}\right)$
This concludes the proof.
Appendix G Proof of Theorem 4
We start with an important result on $A_{\mathbf{v}}(\theta)$.
Lemma 8.
For a large enough $c_{\beta}$, Assumption 3 implies that $A_{\mathbf{v}}(\theta)\succeq(\lambda/2)I$ where $I$ is an identity matrix of appropriate dimension and $\theta$ is an arbitrary policy parameter.
Proof of Lemma 8.
Recall that $A_{\mathbf{v}}(\theta)=\mathbf{E}_{\theta}[\hat{A}_{\mathbf{v}}(z)]$ where $\operatorname{\mathbb{E}}_{\theta}$ denotes expectation over the distribution of $z=(s,a,s^{\prime})$ where $(s,a)\sim\nu^{\pi_{\theta}}$, $s^{\prime}\sim P(\cdot|s,a)$. Hence, for any $\xi=[\eta,\zeta]$, we have
$\begin{split}\xi^{\top}A_{\mathbf{v}}(\theta)\xi&=c_{\beta}\eta^{2}+\eta\zeta^{\top}\operatorname{\mathbb{E}}_{\theta}\left[\phi(s)\right]+\zeta^{\top}\operatorname{\mathbb{E}}_{\theta}\left[\phi(s)\left[\phi(s)-\phi(s^{\prime})\right]^{\top}\right]\zeta\\ &\overset{(a)}{\geq}c_{\beta}\eta^{2}-|\eta|\left\lVert\zeta\right\rVert+\lambda\left\lVert\zeta\right\rVert^{2}\\ &\geq\left\lVert\xi\right\rVert^{2}\left\{\min_{u\in[0,1]}c_{\beta}u-\sqrt{u(1-u)}+\lambda(1-u)\right\}\overset{(b)}{\geq}(\lambda/2)\left\lVert\xi\right\rVert^{2}\end{split}$ (55)
where $(a)$ is a consequence of Assumption 3 and the fact that $\left\lVert\phi(s)\right\rVert\leq 1$, $\forall s\in\mathcal{S}$. Finally, $(b)$ is satisfied when $c_{\beta}\geq\lambda+\sqrt{\frac{1}{\lambda^{2}}-1}$. This concludes the proof of Lemma 8. ∎
Combining Lemma 8 with (51), (52), and (53), we can, therefore, conclude that the following inequalities hold for arbitrary $\theta_{k}$ whenever $c_{\beta}\geq\lambda+\sqrt{\frac{1}{\lambda^{2}}-1}$.
$\dfrac{\lambda}{2}\leq\left\lVert A_{\mathbf{v}}(\theta_{k})\right\rVert\leq\operatorname*{\mathcal{O}}(c_{\beta}),~{}\text{and}~{}\left\lVert b_{\mathbf{v}}(\theta_{k})\right\rVert\leq\operatorname*{\mathcal{O}}(c_{\beta})$ (56)
Utilizing the above result with Lemma 3 and invoking Theorem 2, we arrive at the following.
$\operatorname{\mathbb{E}}_{k}\left[\left\lVert\xi_{k}-\xi_{k}^{*}\right\rVert^{2}\right]\leq\dfrac{1}{H^{2}}\left\lVert\xi_{0}^{k}-\xi_{k}^{*}\right\rVert^{2}+\tilde{\operatorname*{\mathcal{O}}}\left(\dfrac{R_{0}}{H}+R_{1}\right),$
$\operatorname{\mathbb{E}}_{k}\left[\left\lVert\operatorname{\mathbb{E}}_{k}[\xi_{k}]-\xi_{k}^{*}\right\rVert^{2}\right]\leq\dfrac{1}{H^{2}}\left\lVert\xi^{k}_{0}-\xi_{k}^{*}\right\rVert^{2}+\operatorname*{\mathcal{O}}(\bar{R}_{1})+\operatorname*{\mathcal{O}}\left(\dfrac{c_{\beta}^{2}t_{\mathrm{mix}}}{\lambda^{2}H^{2}}\left\{\left\lVert\xi_{0}^{k}-\xi_{k}^{*}\right\rVert^{2}+\operatorname*{\mathcal{O}}\left(R_{0}+R_{1}\right)\right\}\right)$
where the terms $R_{0},R_{1},\bar{R}_{1}$ are defined as follows.
$R_{0}=\tilde{\operatorname*{\mathcal{O}}}\left(\lambda^{-4}c_{\beta}^{4}t_{\mathrm{mix}}+\lambda^{-2}c_{\beta}^{2}t_{\mathrm{mix}}\right)=\tilde{\operatorname*{\mathcal{O}}}\left(\lambda^{-4}c_{\beta}^{4}t_{\mathrm{mix}}\right),$
$R_{1}=\operatorname*{\mathcal{O}}\left(H^{-2}\lambda^{-4}c_{\beta}^{4}t_{\mathrm{mix}}+H^{-2}\lambda^{-2}c_{\beta}^{2}t_{\mathrm{mix}}\right)=\operatorname*{\mathcal{O}}\left(H^{-2}\lambda^{-4}c_{\beta}^{4}t_{\mathrm{mix}}\right)$
$\bar{R}_{1}=\operatorname*{\mathcal{O}}\left(H^{-2}\lambda^{-4}c_{\beta}^{4}t_{\mathrm{mix}}+H^{-2}\lambda^{-2}c_{\beta}^{2}t_{\mathrm{mix}}\right)=\operatorname*{\mathcal{O}}\left(H^{-2}\lambda^{-4}c_{\beta}^{4}t_{\mathrm{mix}}\right)$
Hence, we have the following results.
$\operatorname{\mathbb{E}}_{k}\left[\left\lVert\xi_{k}-\xi_{k}^{*}\right\rVert^{2}\right]\leq\dfrac{1}{H^{2}}\left\lVert\xi_{0}^{k}-\xi_{k}^{*}\right\rVert^{2}+\tilde{\operatorname*{\mathcal{O}}}\left(\dfrac{c_{\beta}^{4}t_{\mathrm{mix}}}{\lambda^{4}H}\right),$
$\left\lVert\operatorname{\mathbb{E}}_{k}[\xi_{k}]-\xi_{k}^{*}\right\rVert^{2}\leq\operatorname*{\mathcal{O}}\left(\dfrac{c_{\beta}^{2}t_{\mathrm{mix}}}{\lambda^{2}H^{2}}\left\lVert\xi_{0}^{k}-\xi_{k}^{*}\right\rVert^{2}\right)+\operatorname*{\mathcal{O}}\left(\dfrac{c_{\beta}^{6}t^{2}_{\mathrm{mix}}}{\lambda^{6}H^{2}}\right)$
This concludes the proof of Theorem 4.
Appendix H Proof of Theorem 1
Recall that the global convergence of any update of form $\theta_{k+1}=\theta_{k}+\alpha\omega_{k}$ can be bounded as
$\begin{split}J^{*}-\frac{1}{K}\sum_{k=0}^{K-1}&\operatorname{\mathbb{E}}[J(\theta_{k})]\leq\sqrt{\epsilon_{\mathrm{bias}}}+\frac{G_{1}}{K}\sum_{k=0}^{K-1}\operatorname{\mathbb{E}}\|(\operatorname{\mathbb{E}}_{k}\left[\omega_{k}\right]-\omega^{*}_{k})\|+\dfrac{\alpha G_{2}}{K}\sum_{k=0}^{K-1}\operatorname{\mathbb{E}}\|\omega_{k}-\omega_{k}^{*}\|^{2}\\ &+\dfrac{\alpha\mu^{-2}}{K}\sum_{k=0}^{K-1}\operatorname{\mathbb{E}}\|\nabla_{\theta}J(\theta_{k})\|^{2}+\frac{1}{\alpha K}\operatorname{\mathbb{E}}_{s\sim d^{\pi^{*}}}[KL(\pi^{*}(\cdot|s)\|\pi_{\theta_{0}}(\cdot|s))].\end{split}$ (57)
We shall now derive a bound for ${\frac{1}{K}\sum_{k=0}^{K-1}\|\nabla_{\theta}J(\theta_{k})\|^{2}}$. Given that the function $J$ is $L$-smooth, we obtain:
$\begin{split}J(\theta_{k+1})&\geq J(\theta_{k})+\left<\nabla_{\theta}J(\theta_{k}),\theta_{k+1}-\theta_{k}\right>-\frac{L}{2}\|\theta_{k+1}-\theta_{k}\|^{2}\\ &=J(\theta_{k})+\alpha\left<\nabla_{\theta}J(\theta_{k}),\omega_{k}\right>-\frac{\alpha^{2}L}{2}\|\omega_{k}\|^{2}\\ &=J(\theta_{k})+\alpha\left<\nabla_{\theta}J(\theta_{k}),\omega_{k}^{*}\right>+\alpha\left<\nabla_{\theta}J(\theta_{k}),\omega_{k}-\omega_{k}^{*}\right>-\frac{\alpha^{2}L}{2}\|\omega_{k}-\omega_{k}^{*}+\omega_{k}^{*}\|^{2}\\ &\overset{(a)}{\geq}J(\theta_{k})+\alpha\left<\nabla_{\theta}J(\theta_{k}),F(\theta_{k})^{-1}\nabla_{\theta}J(\theta_{k})\right>+\alpha\left<\nabla_{\theta}J(\theta_{k}),\omega_{k}-\omega_{k}^{*}\right>\\ &\hskip 28.45274pt-\alpha^{2}L\|\omega_{k}-\omega_{k}^{*}\|^{2}-\alpha^{2}L\|\omega_{k}^{*}\|^{2}\\ &\overset{(b)}{\geq}J(\theta_{k})+\dfrac{\alpha}{G_{1}^{2}}\|\nabla_{\theta}J(\theta_{k})\|^{2}+\alpha\left<\nabla_{\theta}J(\theta_{k}),\omega_{k}-\omega_{k}^{*}\right>-\alpha^{2}L\|\omega_{k}-\omega_{k}^{*}\|^{2}-\alpha^{2}L\|\omega_{k}^{*}\|^{2}\\ &=J(\theta_{k})+\dfrac{\alpha}{2G_{1}^{2}}\|\nabla_{\theta}J(\theta_{k})\|^{2}+\dfrac{\alpha}{2G_{1}^{2}}\left[\|\nabla_{\theta}J(\theta_{k})\|^{2}+2G_{1}^{2}\left<\nabla_{\theta}J(\theta_{k}),\omega_{k}-\omega_{k}^{*}\right>+G_{1}^{4}\|\omega_{k}-\omega_{k}^{*}\|^{2}\right]\\ &\hskip 28.45274pt-\left(\dfrac{\alpha G_{1}^{2}}{2}+\alpha^{2}L\right)\|\omega_{k}-\omega_{k}^{*}\|^{2}-\alpha^{2}L\|\omega_{k}^{*}\|^{2}\\ &=J(\theta_{k})+\dfrac{\alpha}{2G_{1}^{2}}\|\nabla_{\theta}J(\theta_{k})\|^{2}+\dfrac{\alpha}{2G_{1}^{2}}\|\nabla_{\theta}J(\theta_{k})+G_{1}^{2}(\omega_{k}-\omega_{k}^{*})\|^{2}-\left(\dfrac{\alpha G_{1}^{2}}{2}+\alpha^{2}L\right)\|\omega_{k}-\omega_{k}^{*}\|^{2}\\ &\hskip 28.45274pt-\alpha^{2}L\|\omega_{k}^{*}\|^{2}\\ &\geq J(\theta_{k})+\dfrac{\alpha}{2G_{1}^{2}}\|\nabla_{\theta}J(\theta_{k})\|^{2}-\left(\dfrac{\alpha G_{1}^{2}}{2}+\alpha^{2}L\right)\|\omega_{k}-\omega_{k}^{*}\|^{2}-\alpha^{2}L\|F(\theta_{k})^{-1}\nabla_{\theta}J(\theta_{k})\|^{2}\\ &\overset{(c)}{\geq}J(\theta_{k})+\left(\dfrac{\alpha}{2G_{1}^{2}}-\dfrac{\alpha^{2}L}{\mu^{2}}\right)\|\nabla_{\theta}J(\theta_{k})\|^{2}-\left(\dfrac{\alpha G_{1}^{2}}{2}+\alpha^{2}L\right)\|\omega_{k}-\omega_{k}^{*}\|^{2}\end{split}$ (58)
where $(a)$ use the Cauchy-Schwarz inequality and the definition that $\omega_{k}^{*}=F(\theta_{k})^{-1}\nabla_{\theta}J(\theta_{k})$. Relations $(b)$, and $(c)$ are consequences of Assumption 5(a) and 6 respectively. Summing the above inequality over $k\in\{0,\cdots,K-1\}$, rearranging the terms and substituting $\alpha=\frac{\mu^{2}}{4G_{1}^{2}L}$, we obtain
$\begin{split}\dfrac{\mu^{2}}{16G_{1}^{4}L}\left(\dfrac{1}{K}\sum_{k=0}^{K-1}\|\nabla_{\theta}J(\theta_{k})\|^{2}\right)&\leq\dfrac{J(\theta_{K})-J(\theta_{0})}{K}+\left(\dfrac{\mu^{2}}{8L}+\dfrac{\mu^{4}}{16G_{1}^{4}L}\right)\left(\dfrac{1}{K}\sum_{k=0}^{K-1}\|\omega_{k}-\omega_{k}^{*}\|^{2}\right)\\ &\overset{(a)}{\leq}\dfrac{2}{K}+\left(\dfrac{\mu^{2}}{8L}+\dfrac{\mu^{4}}{16G_{1}^{4}L}\right)\left(\dfrac{1}{K}\sum_{k=0}^{K-1}\|\omega_{k}-\omega_{k}^{*}\|^{2}\right)\end{split}$ (59)
where $(a)$ uses the fact that $J(\cdot)$ is absolutely bounded above by $1$. Inequality (59) can be simplified as follows.
$\begin{split}\dfrac{\mu^{-2}}{K}\left(\sum_{k=0}^{K-1}\|\nabla_{\theta}J(\theta_{k})\|^{2}\right)&\leq\dfrac{32LG_{1}^{4}}{\mu^{4}K}+\left(\dfrac{2G_{1}^{4}}{\mu^{2}}+1\right)\left(\dfrac{1}{K}\sum_{k=0}^{K-1}\|\omega_{k}-\omega_{k}^{*}\|^{2}\right)\end{split}$ (60)
Now all that is left is to bound $\mathbb{E}\left[\|\omega_{k}-\omega_{k}^{*}\|^{2}\right]$ and $\|\mathbb{E}_{k}[\omega_{k}]-\omega_{k}^{*}\|$. Assume $\xi^{k}_{0}=0$, $\forall k$. From Theorem 4, we have
$\operatorname{\mathbb{E}}_{k}\left[\left\lVert\xi_{k}-\xi_{k}^{*}\right\rVert^{2}\right]\leq\dfrac{1}{H^{2}}\left\lVert\xi_{k}^{*}\right\rVert^{2}+\tilde{\operatorname*{\mathcal{O}}}\left(\dfrac{c_{\beta}^{4}t_{\mathrm{mix}}}{\lambda^{4}H}\right)\overset{(a)}{=}\tilde{\operatorname*{\mathcal{O}}}\left(\dfrac{c_{\beta}^{4}t_{\mathrm{mix}}}{\lambda^{4}H}\right),$ (61)
$\left\lVert\operatorname{\mathbb{E}}_{k}[\xi_{k}]-\xi_{k}^{*}\right\rVert^{2}\leq\operatorname*{\mathcal{O}}\left(\dfrac{c_{\beta}^{2}t_{\mathrm{mix}}}{\lambda^{2}H^{2}}\left\lVert\xi_{k}^{*}\right\rVert^{2}\right)+\operatorname*{\mathcal{O}}\left(\dfrac{c_{\beta}^{6}t^{2}_{\mathrm{mix}}}{\lambda^{6}H^{2}}\right)\overset{(b)}{=}\operatorname*{\mathcal{O}}\left(\dfrac{c_{\beta}^{6}t^{2}_{\mathrm{mix}}}{\lambda^{6}H^{2}}\right)$ (62)
The relations $(a)$, $(b)$ are due to the fact that $\left\lVert\xi_{k}^{*}\right\rVert^{2}=\left\lVert\left[A_{\mathbf{v}}(\theta_{k})\right]^{-1}b_{\mathbf{v}}(\theta_{k})\right\rVert^{2}\leq\operatorname*{\mathcal{O}}\left(\lambda^{-2}c_{\beta}^{2}\right)$ where the last inequality is a consequence of (56). Assume $\omega_{H}^{k}=0$, $\forall k$. We have the following from Theorem 3.
$\operatorname{\mathbb{E}}_{k}\left[\left\lVert\omega_{k}-\omega_{k}^{*}\right\rVert^{2}\right]$ $\leq\dfrac{1}{H^{2}}\left\lVert\omega_{k}^{*}\right\rVert^{2}+\tilde{\operatorname*{\mathcal{O}}}\left(\dfrac{1}{H}\bigg{\{}\mu^{-4}G_{1}^{6}t_{\mathrm{mix}}^{3}+\mu^{-2}\lambda^{-2}G_{1}^{2}c_{\beta}^{2}t_{\mathrm{mix}}\bigg{\}}\right)+\mu^{-2}G_{1}^{2}\tilde{\operatorname*{\mathcal{O}}}\left(\operatorname{\mathbb{E}}_{k}\left[\left\lVert\xi_{k}-\xi_{k}^{*}\right\rVert^{2}\right]+\epsilon_{\mathrm{app}}\right)$
$\overset{(a)}{\leq}\operatorname*{\mathcal{O}}\left(\dfrac{G_{1}^{2}t^{2}_{\mathrm{mix}}}{\mu^{2}H^{2}}\right)+\tilde{\operatorname*{\mathcal{O}}}\left(\dfrac{1}{H}\bigg{\{}\mu^{-4}G_{1}^{6}t_{\mathrm{mix}}^{3}+\mu^{-2}\lambda^{-2}G_{1}^{2}c_{\beta}^{2}t_{\mathrm{mix}}\bigg{\}}\right)+\dfrac{G_{1}^{2}}{\mu^{2}}\tilde{\operatorname*{\mathcal{O}}}\left(\dfrac{c_{\beta}^{4}t_{\mathrm{mix}}}{\lambda^{4}H}+\epsilon_{\mathrm{app}}\right)$
$\overset{(b)}{\leq}\tilde{\operatorname*{\mathcal{O}}}\left(\dfrac{1}{H}\bigg{\{}\mu^{-4}G_{1}^{6}t_{\mathrm{mix}}^{3}+\mu^{-2}\lambda^{-4}G_{1}^{2}c_{\beta}^{4}t_{\mathrm{mix}}\bigg{\}}\right)+\dfrac{G_{1}^{2}}{\mu^{2}}\operatorname*{\mathcal{O}}\left(\epsilon_{\mathrm{app}}\right)$ (63)
Inequality $(a)$ utilizes the fact that $\left\lVert\omega_{k}^{*}\right\rVert^{2}=\left\lVert F(\theta_{k})^{\dagger}\nabla_{\theta}J(\theta_{k})\right\rVert^{2}\leq\operatorname*{\mathcal{O}}\left(\mu^{-2}G_{1}^{2}t_{\mathrm{mix}}^{2}\right)$ where the last inequality follows from Assumption 5, 6, and Lemma 7. We also apply (61) to prove $(a)$ whereas $(b)$ is established by retaining only the dominant terms. Theorem 3 also states that
$\left\lVert\operatorname{\mathbb{E}}_{k}[\omega_{k}]-\omega_{k}^{*}\right\rVert^{2}\leq\operatorname*{\mathcal{O}}\left(\mu^{-2}G_{1}^{2}\left\lVert\operatorname{\mathbb{E}}_{k}[\xi_{k}]-\xi_{k}^{*}\right\rVert^{2}+\mu^{-2}G_{1}^{2}\epsilon_{\mathrm{app}}\right)$
$\hskip 28.45274pt+\operatorname*{\mathcal{O}}\left(\dfrac{G_{1}^{4}t_{\mathrm{mix}}}{\mu^{2}H^{2}}\left\{\left\lVert\omega_{k}^{*}\right\rVert^{2}+\mu^{-2}G_{1}^{2}t_{\mathrm{mix}}\operatorname{\mathbb{E}}_{k}\left[\left\lVert\xi_{k}-\xi_{k}^{*}\right\rVert^{2}\right]+\mu^{-4}G_{1}^{6}t_{\mathrm{mix}}^{3}+\mu^{-2}\lambda^{-2}G_{1}^{2}c_{\beta}^{2}t_{\mathrm{mix}}\right\}\right)$
$\overset{(a)}{\leq}\tilde{\operatorname*{\mathcal{O}}}\left(\dfrac{G_{1}^{2}c_{\beta}^{6}t^{2}_{\mathrm{mix}}}{\mu^{2}\lambda^{6}H^{2}}+\dfrac{G_{1}^{2}}{\mu^{2}}\epsilon_{\mathrm{app}}\right)+\tilde{\operatorname*{\mathcal{O}}}\left(\dfrac{1}{H^{2}}\left\{\mu^{-4}G_{1}^{6}t_{\mathrm{mix}}^{3}+\dfrac{G_{1}^{6}c_{\beta}^{4}t^{3}_{\mathrm{mix}}}{\mu^{4}\lambda^{4}H}+\mu^{-6}G_{1}^{10}t_{\mathrm{mix}}^{4}+\mu^{-4}\lambda^{-2}G_{1}^{6}c_{\beta}^{2}t^{2}_{\mathrm{mix}}\right\}\right)$
$\overset{(b)}{\leq}\tilde{\operatorname*{\mathcal{O}}}\left(\dfrac{1}{H^{2}}\bigg{\{}\mu^{-6}G_{1}^{10}t_{\mathrm{mix}}^{4}+\mu^{-2}\lambda^{-6}G_{1}^{2}c_{\beta}^{6}t^{2}_{\mathrm{mix}}+\mu^{-4}\lambda^{-2}G_{1}^{6}c_{\beta}^{2}t^{2}_{\mathrm{mix}}\bigg{\}}\right)+\dfrac{G_{1}^{2}}{\mu^{2}}\operatorname*{\mathcal{O}}\left(\epsilon_{\mathrm{app}}\right)$ (64)
where $(a)$ is a consequence of (61), (62), and the upper bound $\left\lVert\omega_{k}^{*}\right\rVert^{2}\leq\operatorname*{\mathcal{O}}\left(\mu^{-2}G_{1}^{2}t_{\mathrm{mix}}^{2}\right)$ derived earlier. Inequality $(b)$ retains only the dominant terms. Combining (57), (60), (H), and (H), we arrive at the following.
$J^{*}-\frac{1}{K}\sum_{k=0}^{K-1}$ $\operatorname{\mathbb{E}}[J(\theta_{k})]\leq\sqrt{\epsilon_{\mathrm{bias}}}+\tilde{\operatorname*{\mathcal{O}}}\left(\dfrac{1}{H}\bigg{\{}\mu^{-3}G_{1}^{6}t_{\mathrm{mix}}^{2}+\mu^{-1}\lambda^{-3}G_{1}^{2}c^{3}_{\beta}t_{\mathrm{mix}}+\mu^{-2}\lambda^{-1}G_{1}^{4}c_{\beta}t_{\mathrm{mix}}\bigg{\}}\right)+\dfrac{G_{1}^{2}}{\mu}\operatorname*{\mathcal{O}}\left(\sqrt{\epsilon_{\mathrm{app}}}\right)$
$+\dfrac{1}{L}\left(G_{1}^{2}+\dfrac{\mu^{2}G_{2}}{G_{1}^{2}}\right)\left[\tilde{\operatorname*{\mathcal{O}}}\left(\dfrac{1}{H}\bigg{\{}\mu^{-4}G_{1}^{6}t_{\mathrm{mix}}^{3}+\mu^{-2}\lambda^{-4}G_{1}^{2}c_{\beta}^{4}t_{\mathrm{mix}}\bigg{\}}\right)+\dfrac{G_{1}^{2}}{\mu^{2}}\operatorname*{\mathcal{O}}\left(\epsilon_{\mathrm{app}}\right)\right]+\operatorname*{\mathcal{O}}\left(\dfrac{G_{1}^{2}L}{\mu^{2}K}\right)$ (65)
We get the desired result by substituting the values of $H,K$ as stated in the theorem. We want to emphasize that the $G_{1}^{2}$ factor with the $\sqrt{\epsilon_{\mathrm{app}}}$ term is a standard component in actor-critic results with a linear critic [Suttle et al., 2023, Patel et al., 2024]. However, this factor is often not explicitly mentioned in previous works, whereas we have included it here.